### v1

# Adaptive Cohen’s Class Time-Frequency Distribution

Abstract
The fixed kernel function-based Cohen’s class time-frequency distributions (CCTFDs) allow flexibility in denoising for some specific polluted signals. Due to the limitation of fixed kernel functions, however, from the view point of filtering they fail to automatically adjust the response according to the change of signal to adapt to different signal characteristics. In this letter, we integrate Wiener filter principle and the time-frequency filtering mechanism of CCTFD to design the least-squares adaptive filter method in the Wigner-Ville distribution (WVD) domain, giving birth to the least-squares adaptive filter-based CCTFD whose kernel function can be adjusted with the input signal automatically to achieve the minimum mean-square error denoising in the WVD domain. Some examples are also carried out to demonstrate that the proposed adaptive CCTFD outperforms some state-of-the-arts in noise suppression.
Index Terms:
Cohen’s class time-frequency distribution, convolution, least-squares adaptive filter, mean-square error, power spectral density.
I Introduction
Cohen’s class time-frequency distribution (CCTFD) [1], also known as the bi-linear kernel function time-frequency distribution, is one of the most representative time-frequency analysis tools of the conventional time-frequency distributions. It includes particular cases the Wigner-Ville distribution (WVD) [2], the Choi-Williams distribution [3], the Kirkwood-Rihaczek distribution [4], the Born-Jordan distribution [5], the Zhao-Atlas-Marks distribution [6], the Margenau-Hill distribution [7], and the Page distribution [8]. Indeed, it can be regarded as a unified bi-linear time-frequency distribution and has found many applications in seismic exploration, electronic countermeasures, deep-sea detection, spectral imaging, and ultrasonic inspection [9, 10, 11, 12, 13, 14].
From the view point of signal processing, the integral form of CCTFD can be rewritten as the conventional convolution of the WVD and the kernel function. Namely, the CCTFD is a result of smoothing the WVD by using the kernel function. The main purpose the smoothing filter is not just only to suppress cross-terms of the WVD, but to achieve non-stationary signal denoising in the WVD domain, and accordingly extract the target signal from background noises or eliminate the impact of noises through the unique reconstruction property of WVD.
There are a number of attempts to use the filtering instinct of CCTFD to remove noises, giving birth to many useful and effective kernel functions of which the CCTFD works well [15, 16, 17, 18, 19]. These filtering methods employ fixed kernel functions in dealing with specific input signals, which limits the scope of applications severely. To overcome this shortcoming, it seems feasible to explore the adaptive CCTFD whose kernel function can be adjusted with the input signal automatically.
Adaptive filter is a kind of filter that is able to automatically adjust its parameters or response according to the change of signal to adapt to different signal characteristics [20]. The most celebrated result in this field is Wiener’s result [21] which applies the minimum mean-square error (MSE) criterion to design the least-squares adaptive filter method. The Wiener filter has been widely used in addressing practical issues encountered frequently in radar, communications, sonar, biomedicine, and vibration engineering [22, 23, 24, 25, 26, 27, 28].
The core idea of this letter is to integrate Wiener filter principle and the time-frequency filtering mechanism of CCTFD to investigate the convolution type of CCTFD time-frequency analysis method. The main purpose of this letter is to design the least-squares adaptive filter method in the WVD domain, disclose the influence mechanism of the kernel function on the effect of denoising, and establish the adaptive kernel function of CCTFD with the minimum MSE in the WVD domain. The main contributions of this letter are summarized as follows:
  • This letter obtains the least-squares adaptive filter in the WVD domain.
  • This letter proposes the adaptive CCTFD whose kernel function takes the reversal of the least-squares adaptive filter transfer function in the WVD domain to restore signals from an extreme strong additive noise background.
  • This letter demonstrates the noise suppression superiority of the proposed adaptive CCTFD over the ordinary Wiener filter and some classical fixed kernel function-based CCTFDs.
The remainder of this letter is structured as follows. In Section II, we recall some necessary background and notation on the convolution type of CCTFD. In Section III, we provide the CCTFD-based adaptive filter method for additive noises jamming signals. In Section IV, we introduce some examples to validate the effectiveness, reliability and feasibility of the proposed method. In Section V, we draw a conclusion. All the technical proofs of our theoretical results are relegated to the appendix parts.
To facilitate the understanding and clear representation of the various symbols used throughout the paper, we present Table I, which provides a detailed explanation of each symbol and its corresponding meaning.
TABLE I: Symbol description.
Symbole Description
$\mathrm{T}$ transpose operator
complex conjugate operator
$\ast$ convolution operator
$\left\|\cdot\right\|_{2}$ $L^{2}$-norm operator
$\mathbb{E}(\cdot)$ mathematical expectation operator
$R$ correlation function operator
$\mathrm{C}$ CCTFD operator
$\mathrm{W}$ WVD operator
$\mathcal{F}$ FT operator
$\phi$ kernel function
$\Pi$ FT of the kernel function $\phi$
$H$ adaptive filter in the WVD domain
$\sigma_{\mathrm{MSE}}^{2}$ MSE
$\varepsilon$ PSD
$\mathrm{C}_{g}^{\mathrm{LSAF}}$ the least-squares adaptive filter-based CCTFD for the input signal $g$
II Convolution type of CCTFD
The integral form of CCTFD of the function $f(\boldsymbol{x})\in L^{2}(\mathbb{R}^{N})$ reads [1, 29]
$\mathrm{C}_{f}(\boldsymbol{x},\boldsymbol{w})=$ $\int_{\mathbb{R}^{N}}\int_{\mathbb{R}^{N}}\int_{\mathbb{R}^{N}}f\left(\boldsymbol{y}+\frac{\boldsymbol{\tau}}{2}\right)\overline{f\left(\boldsymbol{y}-\frac{\boldsymbol{\tau}}{2}\right)}$
×$\times\phi(\boldsymbol{\theta},\boldsymbol{\tau})\mathrm{e}^{-2\pi\mathrm{i}\left(\boldsymbol{\theta}\boldsymbol{x}^{\mathrm{T}}+\boldsymbol{\tau}\boldsymbol{w}^{\mathrm{T}}-\boldsymbol{y}\boldsymbol{\theta}^{\mathrm{T}}\right)}\mathrm{d}\boldsymbol{y}\mathrm{d}\boldsymbol{\tau}\mathrm{d}\boldsymbol{\theta},$ (1)
where the superscripts $\mathrm{T}$ and — denote the transpose operator and complex conjugate operator, respectively, and $\phi(\boldsymbol{\theta},\boldsymbol{\tau})$ denotes the kernel function.
Let $\ast$ and $\mathcal{F}$ be the conventional convolution operator and Fourier operator, respectively. Then, Eq. (II) can be rewritten as
$\mathrm{C}_{f}(\boldsymbol{x},\boldsymbol{w})=$ $\left(\mathrm{W}_{f}\ast\Pi\right)(\boldsymbol{x},\boldsymbol{w}),$ (2)
where
$\mathrm{W}_{f}(\boldsymbol{x},\boldsymbol{w})=\int_{\mathbb{R}^{N}}f\left(\boldsymbol{x}+\frac{\boldsymbol{\tau}}{2}\right)\overline{f\left(\boldsymbol{x}-\frac{\boldsymbol{\tau}}{2}\right)}\mathrm{e}^{-2\pi\mathrm{i}\boldsymbol{\tau}\boldsymbol{w}^{\mathrm{T}}}\mathrm{d}\boldsymbol{\tau}$ (3)
denotes the WVD of the function $f(\boldsymbol{x})$, and
$\Pi(\boldsymbol{x},\boldsymbol{w})=$ $\mathcal{F}[\phi](\boldsymbol{x},\boldsymbol{w})$
$=$ $\int_{\mathbb{R}^{N}}\int_{\mathbb{R}^{N}}\phi(\boldsymbol{\theta},\boldsymbol{\tau})\mathrm{e}^{-2\pi\mathrm{i}\left(\boldsymbol{\theta}\boldsymbol{x}^{\mathrm{T}}+\boldsymbol{\tau}\boldsymbol{w}^{\mathrm{T}}\right)}\mathrm{d}\boldsymbol{\theta}\mathrm{d}\boldsymbol{\tau}$ (4)
denotes the Fourier transform (FT) of the kernel function $\phi(\boldsymbol{\theta},\boldsymbol{\tau})$. Eq. (2) indicates that the CCTFD is none other than the convolution of the WVD and the (FT version of) kernel function.
III Least-squares adaptive filter-based CCTFD
For a given noise polluted signal $g(\boldsymbol{x})=f(\boldsymbol{x})+n(\boldsymbol{x})$, where $f(\boldsymbol{x})$ and $n(\boldsymbol{x})$ denote the pure signal and the additive noise, respectively, the common tactic of filters is to restore the pure signal as accurately as possible, namely, find the estimate $\widehat{f}(\boldsymbol{x})$ as close as possible to the ideal $f(\boldsymbol{x})$. Thanks to the convolution nature of CCTFD and the unique reconstruction property of WVD, this is equivalent to design an adaptive filter $H(\boldsymbol{x},\boldsymbol{w})$ in the WVD domain, which can find the estimate
$\mathrm{W}_{\widehat{f}}(\boldsymbol{x},\boldsymbol{w})=\left(\mathrm{W}_{g}\ast H\right)(\boldsymbol{x},\boldsymbol{w})$ (5)
as close as possible to the ideal $\mathrm{W}_{f}(\boldsymbol{x},\boldsymbol{w})$. According to Wiener filter principle, a natural criterion to characterize the estimation accuracy is the MSE criterion
$\sigma_{\mathrm{MSE}}^{2}\overset{\mathrm{def}}{=}\mathbb{E}\left\{\left|\mathrm{W}_{f}(\boldsymbol{x},\boldsymbol{w})-\mathrm{W}_{\widehat{f}}(\boldsymbol{x},\boldsymbol{w})\right|^{2}\right\},$ (6)
where $\mathbb{E}(\cdot)$ denotes the mathematical expectation operator.
For simplicity, let $\boldsymbol{z}=(\boldsymbol{x},\boldsymbol{w})\in\mathbb{R}^{2N}$, then Eqs. (5) and (6) become
$\mathrm{W}_{\widehat{f}}(\boldsymbol{z})=\left(\mathrm{W}_{g}\ast H\right)(\boldsymbol{z})=\int_{\mathbb{R}^{2N}}\mathrm{W}_{g}(\boldsymbol{k})H(\boldsymbol{z}-\boldsymbol{k})\mathrm{d}\boldsymbol{k}$ (7)
and
$\sigma_{\mathrm{MSE}}^{2}=\mathbb{E}\left\{\left|\mathrm{W}_{f}(\boldsymbol{z})-\mathrm{W}_{\widehat{f}}(\boldsymbol{z})\right|^{2}\right\},$ (8)
respectively. Now, our goal is to design an adaptive optimal filter $H_{\mathrm{opt}}(\boldsymbol{z})$ in the WVD domain to minimize the MSE given by Eq. (8), or equivalently,
$H_{\mathrm{opt}}(\boldsymbol{z})=\mathop{\arg\min}\limits_{H(\boldsymbol{z})}\sigma_{\mathrm{MSE}}^{2}.$ (9)
By using the orthogonal principle [30], the stationary assumption and the conventional convolution and correlation theorems to establish, simplify and solve the Wiener-Hopf equation, respectively, the least-squares adaptive filter transfer function in the WVD domain reads
$\mathcal{F}\left[H_{\mathrm{opt}}\right](\boldsymbol{u})=\frac{\varepsilon_{\mathrm{W}_{f},\mathrm{W}_{g}}(\boldsymbol{u})}{\varepsilon_{\mathrm{W}_{g}}(\boldsymbol{u})},$ (10)
where $\varepsilon_{\mathrm{W}_{f},\mathrm{W}_{g}}(\boldsymbol{u})=\mathcal{F}\left[\mathrm{W}_{f}\right](\boldsymbol{u})\overline{\mathcal{F}\left[\mathrm{W}_{g}\right](\boldsymbol{u})}$ denotes the power spectral density (PSD) of $\mathrm{W}_{f}(\boldsymbol{z})$ and $\mathrm{W}_{g}(\boldsymbol{z})$, and $\varepsilon_{\mathrm{W}_{g}}(\boldsymbol{u})=\left|\mathcal{F}\left[\mathrm{W}_{g}\right](\boldsymbol{u})\right|^{2}$ denotes the PSD of $\mathrm{W}_{g}(\boldsymbol{z})$. Taking the inverse FT on both sides of Eq. (10) yields the least-squares adaptive filter in the WVD domain
$H_{\mathrm{opt}}(\boldsymbol{z})=\int_{\mathbb{R}^{2N}}\frac{\varepsilon_{\mathrm{W}_{f},\mathrm{W}_{g}}(\boldsymbol{u})}{\varepsilon_{\mathrm{W}_{g}}(\boldsymbol{u})}\mathrm{e}^{2\pi\mathrm{i}\boldsymbol{u}\boldsymbol{z}^{\mathrm{T}}}\mathrm{d}\boldsymbol{u}.$ (11)
See Appendix A for the detailed derivation of Eq. (10). Correspondingly, the minimum MSE can be reduced to zero, i.e.
$\mathop{\min}\limits_{H(\boldsymbol{z})}\sigma_{\mathrm{MSE}}^{2}=0,$ (12)
See Appendix B for the detailed derivation of Eq. (12).
From Eq. (II), it derives that the adaptive optimal kernel function takes the reversal of the least-squares adaptive filter transfer function in the WVD domain, i.e.,
$\phi_{\mathrm{opt}}(\boldsymbol{\theta},\boldsymbol{\tau})=\mathcal{F}\left[H_{\mathrm{opt}}\right](-\boldsymbol{\theta},-\boldsymbol{\tau})=\frac{\varepsilon_{\mathrm{W}_{f},\mathrm{W}_{g}}(-\boldsymbol{\theta},-\boldsymbol{\tau})}{\varepsilon_{\mathrm{W}_{g}}(-\boldsymbol{\theta},-\boldsymbol{\tau})}.$ (13)
Substituting Eq. (13) into Eq. (II) gives the least-squares adaptive filter-based CCTFD
$\mathrm{C}_{g}^{\mathrm{LSAF}}(\boldsymbol{x},\boldsymbol{w})$
$=$ $\int_{\mathbb{R}^{N}}\int_{\mathbb{R}^{N}}\int_{\mathbb{R}^{N}}g\left(\boldsymbol{y}+\frac{\boldsymbol{\tau}}{2}\right)\overline{g\left(\boldsymbol{y}-\frac{\boldsymbol{\tau}}{2}\right)}$
×$\times\frac{\varepsilon_{\mathrm{W}_{f},\mathrm{W}_{g}}(-\boldsymbol{\theta},-\boldsymbol{\tau})}{\varepsilon_{\mathrm{W}_{g}}(-\boldsymbol{\theta},-\boldsymbol{\tau})}\mathrm{e}^{-2\pi\mathrm{i}\left(\boldsymbol{\theta}\boldsymbol{x}^{\mathrm{T}}+\boldsymbol{\tau}\boldsymbol{w}^{\mathrm{T}}-\boldsymbol{y}\boldsymbol{\theta}^{\mathrm{T}}\right)}\mathrm{d}\boldsymbol{y}\mathrm{d}\boldsymbol{\tau}\mathrm{d}\boldsymbol{\theta}.$ (14)
It is obvious that the transformation is designed to adapt based on the input signals $f$ and $g$. This adaptability is achieved through the use of different CCTFD kernel functions, which are selected and optimized according to the characteristics of the input signals. As a result, the adaptive filter dynamically adjusts to varying signal conditions, providing a tailored and efficient filtering process that enhances the clarity and accuracy of the time-frequency representation.
IV Numerical experiments
This section performs four examples to verify the correctness and effectiveness of the least-squares adaptive filter method in the WVD domain. It also compares the denoising effect of the proposed adaptive CCTFD with that of the ordinary Wiener filter and some classical fixed kernel function-based CCTFDs.
In simulations, the fixed kernel functions are chosen as $\phi(\boldsymbol{\theta},\boldsymbol{\tau})=\cos\left(\frac{\boldsymbol{\theta}\boldsymbol{\tau}^{\mathrm{T}}}{2}\right)$, $\phi(\boldsymbol{\theta},\boldsymbol{\tau})=\mathrm{e}^{\mathrm{i}\frac{\boldsymbol{\theta}\boldsymbol{\tau}^{\mathrm{T}}}{2}}$, $\phi(\boldsymbol{\theta},\boldsymbol{\tau})=\frac{\sin\left(\frac{\boldsymbol{\theta}\boldsymbol{\tau}^{\mathrm{T}}}{2}\right)}{\frac{\boldsymbol{\theta}\boldsymbol{\tau}^{\mathrm{T}}}{2}}$ and $\phi(\boldsymbol{\theta},\boldsymbol{\tau})=\mathrm{e}^{\mathrm{i}\boldsymbol{\theta}\left\|\boldsymbol{\tau}\right\|_{1}}$, (here $\left\|\cdot\right\|_{1}$ denotes the $1$-norm for vectors), corresponding to the Margenau-Hill distribution, the Kirkwood-Rihaczek distribution, the Born-Jordan distribution and the Page distribution, respectively. For simplicity, we call the filtering methods using fixed kernel functions $\cos\left(\frac{\boldsymbol{\theta}\boldsymbol{\tau}^{\mathrm{T}}}{2}\right)$, $\mathrm{e}^{\mathrm{i}\frac{\boldsymbol{\theta}\boldsymbol{\tau}^{\mathrm{T}}}{2}}$, $\frac{\sin\left(\frac{\boldsymbol{\theta}\boldsymbol{\tau}^{\mathrm{T}}}{2}\right)}{\frac{\boldsymbol{\theta}\boldsymbol{\tau}^{\mathrm{T}}}{2}}$ and $\mathrm{e}^{\mathrm{i}\boldsymbol{\theta}\left\|\boldsymbol{\tau}\right\|_{1}}$ as Margenau-Hill, Kirkwood-Rihaczek, Born-Jordan and Page, respectively.
Example 1 (Linear frequency-modulated (LFM) signal): The polluted signal is selected for the LFM signal $\mathrm{e}^{2\pi\mathrm{i}\left(\boldsymbol{x}+\frac{\boldsymbol{x}^{2}}{2}\right)}$.
Example 2 (Gaussian enveloped LFM (GELFM) signal): The polluted signal is selected for the GELFM signal $\mathrm{e}^{-\frac{(\boldsymbol{x}+1)^{2}}{8}}\mathrm{e}^{2\pi\mathrm{i}\boldsymbol{x}^{2}}$.
Example 3 (Quadratic frequency-modulated (QFM) signal): The polluted signal is selected for the QFM signal $\mathrm{e}^{2\pi\mathrm{i}\left(-3\boldsymbol{x}+\frac{\boldsymbol{x}^{2}}{2}+\frac{\boldsymbol{x}^{3}}{4}\right)}$.
Example 4 (Sinusoidal frequency-modulated (SFM) signal): The polluted signal is selected for the SFM signal $\mathrm{e}^{\mathrm{i}\left[1.3\pi\boldsymbol{x}+2\sin(0.6\pi\boldsymbol{x})\right]}$.
In examples 1–4, additive white Gaussian noise is introduced to the signals, with the SNR ranging from $-5\mathrm{dB}$ to $5\mathrm{dB}$ and the observation interval is set to $[-5\mathrm{s},5\mathrm{s}]$. The sampling frequencies are set as follows: $80\mathrm{Hz}$ for example 1, $100\mathrm{Hz}$ for example 2, $150\mathrm{Hz}$ for example 3 and $175\mathrm{Hz}$ for example 4.
Figs. 1(a) and (b) plot respectively the SNR-MSE (logarithm base $10$) and SNR-peak SNR (PSNR) (average of real and imaginary parts) line charts of the estimated LFM siganl using six filtering methods including Margenau-Hill, Kirkwood-Rihaczek, Born-Jordan, Page, Wiener filter and the proposed adaptive CCTFD. Figs. 2(a) and (b) plot respectively the SNR-MSE and SNR-PSNR line charts of the estimate GELFM signals using these six filtering methods. Figs. 3(a) and (b) plot respectively the SNR-MSE and SNR-PSNR line charts of the estimate QFM signals using these six filtering methods. Figs. 4(a) and (b) plot respectively the SNR-MSE and SNR-PSNR line charts of the estimate SFM signals using these six filtering methods. It can be seen that the proposed adaptive CCTFD achieves better noise suppression performance than Margenau-Hill, Kirkwood-Rihaczek, Born-Jordan, Page, and Wiener filters under different SNR levels.
Figure 1: Performance evaluation of six denoising methods for the LFM signal at SNR levels from $-5\mathrm{dB}$ to $5\mathrm{dB}$ under white Gaussian noise: (a) MSE performance of the estimated LFM signals; (b) PSNR performance of the estimated LFM signals.
Figure 2: Performance evaluation of six denoising methods for the GELFM signal at SNR levels from $-5\mathrm{dB}$ to $5\mathrm{dB}$ under white Gaussian noise: (a) MSE performance of the estimated GELFM signals; (b) PSNR performance of the estimated GELFM signals.
Figure 3: Performance evaluation of six denoising methods for the QFM signal at SNR levels from $-5\mathrm{dB}$ to $5\mathrm{dB}$ under white Gaussian noise: (a) MSE performance of the estimated QFM signals; (b) PSNR performance of the estimated QFM signals.
Figure 4: Performance evaluation of six denoising methods for the SFM signal at SNR levels from $-5\mathrm{dB}$ to $5\mathrm{dB}$ under white Gaussian noise: (a) MSE performance of the estimated SFM signals; (b) PSNR performance of the estimated SFM signals.
Additionally, we investigated the denoising performance of the six methods under conditions of different colored noise. The four signals were subjected to various types of colored noise: pink noise, blue noise, and red noise. The observation interval is set to $[-5\mathrm{s},5\mathrm{s}]$, and the sampling frequencies are set as follows: $30\mathrm{Hz}$ for example 1, $50\mathrm{Hz}$ for example 2, $150\mathrm{Hz}$ for examples 3 and 4. Table II presents the MSE and PSNR values of the output signals through the six filters under the influence of three types of colored noise. It can be seen that for blue noise, our proposed adaptive CCTFD significantly outperforms the other methods. For pink and red noise, our method is superior to the others for LFM and GELFM signals (except for the PSNR value of red noise in the GELFM signal, which is slightly lower than that of the Wiener filter). For pink and red noise, the Wiener filter performs better than ours for QFM and SFM signals. However, overall, our method consistently outperforms the fixed kernel function filtering methods.
TABLE II: The MSE and PSNR of six filtering methods under different colored noise conditions.
Example LFM GELFM QFM SFM
MSE (Mean-Square Error)
Pink noise Margenau-Hill $-0.0016$ $-0.1592$ $-0.0513$ $-0.1319$
Kirkwood-Rihaczek $0.0676$ $-0.0050$ $-0.0330$ $-0.1200$
Born-Jordan $-0.1582$ $-0.3053$ $-0.1123$ $-0.2134$
Page $0.0996$ $-0.0360$ $-0.0201$ $-0.0582$
Wiener filter $-0.4985$ $-0.6911$ $\mathbf{-0.4699}$ $\mathbf{-0.4268}$
Adaptive CCTFD $\mathbf{-1.5318}$ $\mathbf{-0.9807}$ $-0.3169$ $-0.3369$
Blue noise Margenau-Hill $0.1437$ $-0.4563$ $-0.0637$ $-0.3730$
Kirkwood-Rihaczek $0.3580$ $-0.3711$ $0.0063$ $-0.2404$
Born-Jordan $-0.1057$ $-0.5265$ $-0.2016$ $-0.9358$
Page $0.3281$ $-0.3355$ $0.0239$ $0.0682$
Wiener filter $-0.4676$ $-0.6632$ $-0.5208$ $-1.0831$
Adaptive CCTFD $\mathbf{-1.2999}$ $\mathbf{-0.9581}$ $\mathbf{-0.7917}$ $\mathbf{-1.6128}$
Red noise Margenau-Hill $0.0655$ $-0.3004$ $-0.0624$ $-0.2233$
Kirkwood-Rihaczek $0.1243$ $-0.2567$ $-0.0475$ $-0.2145$
Born-Jordan $-0.0907$ $-0.3810$ $-0.1369$ $-0.3623$
Page $0.1720$ $-0.2790$ $-0.0082$ $-0.0851$
Wiener filter $-0.8039$ $-1.2798$ $\mathbf{-0.9489}$ $\mathbf{-0.7804}$
Adaptive CCTFD $\mathbf{-1.7448}$ $\mathbf{-1.3028}$ $-0.4032$ $-0.3919$
PSNR (Peak Signal-to-Noise Ratio)
Pink noise Margenau-Hill $3.0323$ $5.2766$ $3.5237$ $4.3306$
Kirkwood-Rihaczek $2.3400$ $3.7410$ $3.3431$ $4.2110$
Born-Jordan $4.6190$ $7.0716$ $4.1339$ $5.1494$
Page $2.0433$ $3.7370$ $3.2109$ $3.5963$
Wiener filter $8.8314$ $10.3192$ $\mathbf{8.0556}$ $\mathbf{7.8716}$
Adaptive CCTFD $\mathbf{18.5390}$ $\mathbf{12.8503}$ $6.1978$ $6.5799$
Blue noise Margenau-Hill $1.5865$ $7.6228$ $3.6468$ $7.2377$
Kirkwood-Rihaczek $-0.5454$ $6.8108$ $2.9469$ $5.4278$
Born-Jordan $4.1557$ $8.3757$ $5.0342$ $13.0179$
Page $-0.2493$ $6.4764$ $2.7708$ $2.3297$
Wiener filter $8.1082$ $9.8937$ $8.6899$ $14.4000$
Adaptive CCTFD $\mathbf{16.0114}$ $\mathbf{12.6000}$ $\mathbf{10.9354}$ $\mathbf{19.6753}$
Red noise Margenau-Hill $2.3888$ $6.3998$ $3.6360$ $5.2562$
Kirkwood-Rihaczek $1.7774$ $5.9456$ $3.4881$ $5.1559$
Born-Jordan $4.0111$ $7.4167$ $4.3805$ $6.6348$
Page $1.3606$ $5.9863$ $3.0925$ $3.8628$
Wiener filter $11.8832$ $\mathbf{16.7052}$ $\mathbf{13.1156}$ $\mathbf{11.9071}$
Adaptive CCTFD $\mathbf{20.4654}$ $16.1495$ $7.0496$ $7.2056$
V Conclusion
The convolution type of CCTFD time-frequency analysis method suitable for the adaptive filter denoising for additive noises jamming signals under the condition of low SNR was established, giving rise to the adaptive kernel function which minimizes the MSE in the WVD domain. The proposed adaptive CCTFD can automatically adjust its kernel function according to the change of signal to adapt to different signal characteristics. It turns out that its denoising effect is better than that of some classical fixed kernel function-based CCTFDs, including the Margenau-Hill distribution, the Kirkwood-Rihaczek distribution, the Born-Jordan distribution, and the Page distribution. Under white Gaussian noise, our method outperforms the ordinary Wiener filter, while under colored noise, it performs comparably to the Wiener filter.
Appendix
In this section, we prove all the theoretical results in the letter.
V-A Proof of Eq. (10)
By using the orthogonal principle
$\mathbb{E}\left\{\left[\mathrm{W}_{f}(\boldsymbol{z})-\left(\mathrm{W}_{g}\ast H_{\mathrm{opt}}\right)(\boldsymbol{z})\right]\overline{\mathrm{W}_{g}(\boldsymbol{z^{\prime}})}\right\}=0,\boldsymbol{z^{\prime}}\in\mathbb{R}^{2N},$ (A.1)
we establish the Wiener-Hopf equation
$R_{\mathrm{W}_{f},\mathrm{W}_{g}}(\boldsymbol{z},\boldsymbol{z^{\prime}})-\int_{\mathbb{R}^{2N}}R_{\mathrm{W}_{g}}(\boldsymbol{k},\boldsymbol{z^{\prime}})H_{\mathrm{opt}}(\boldsymbol{z}-\boldsymbol{k})\mathrm{d}\boldsymbol{k}=0,$ (A.2)
where $R_{\mathrm{W}_{f},\mathrm{W}_{g}}$ denotes the cross-correlation function between $\mathrm{W}_{f}$ and $\mathrm{W}_{g}$, and $R_{\mathrm{W}_{g}}$ denotes the auto-correlation function of $\mathrm{W}_{g}$. In general, we can obtain $H_{\mathrm{opt}}$ by solving Eq. (A.2) numerically. Particularly, if $\mathrm{W}_{f}$ and $\mathrm{W}_{g}$ are stationary, Eq. (A.2) simplifies to
$R_{\mathrm{W}_{f},\mathrm{W}_{g}}(\boldsymbol{z}-\boldsymbol{z^{\prime}})-\int_{\mathbb{R}^{2N}}R_{\mathrm{W}_{g}}(\boldsymbol{k}-\boldsymbol{z^{\prime}})H_{\mathrm{opt}}(\boldsymbol{z}-\boldsymbol{k})\mathrm{d}\boldsymbol{k}=0.$ (A.3)
Taking the change of variables $\boldsymbol{z}-\boldsymbol{z^{\prime}}=\boldsymbol{p}$ and $\boldsymbol{z}-\boldsymbol{k}=\boldsymbol{q}$ yields
$R_{\mathrm{W}_{f},\mathrm{W}_{g}}(\boldsymbol{p})-\int_{\mathbb{R}^{2N}}R_{\mathrm{W}_{g}}(\boldsymbol{p}-\boldsymbol{q})H_{\mathrm{opt}}(\boldsymbol{q})\mathrm{d}\boldsymbol{q}=0.$ (A.4)
Thanks to the conventional convolution and correlation theorems, we solve Eq. (A.4) to obtain
$\mathcal{F}\left[\mathrm{W}_{f}\right](\boldsymbol{u})\overline{\mathcal{F}\left[\mathrm{W}_{g}\right](\boldsymbol{u})}=\mathcal{F}\left[H_{\mathrm{opt}}\right](\boldsymbol{u})\left|\mathcal{F}\left[\mathrm{W}_{g}\right](\boldsymbol{u})\right|^{2},$ (A.5)
[…]
V-B Proof of Eq. (12)
The minimize MSE takes
$\mathop{\min}\limits_{H(\boldsymbol{z})}\sigma_{\mathrm{MSE}}^{2}=\mathbb{E}\left\{\left[\mathrm{W}_{f}(\boldsymbol{z})-\left(\mathrm{W}_{g}\ast H_{\mathrm{opt}}\right)(\boldsymbol{z})\right]\overline{\mathrm{W}_{f}(\boldsymbol{z})}\right\}.$ (B.1)
Similar to Eqs. (A.2)–(A.4), we have
$\mathop{\min}\limits_{H(\boldsymbol{z})}\sigma_{\mathrm{MSE}}^{2}=R_{\mathrm{W}_{f}}(\boldsymbol{0})-\int_{\mathbb{R}^{2N}}R_{\mathrm{W}_{g},\mathrm{W}_{f}}(-\boldsymbol{k})H_{\mathrm{opt}}(\boldsymbol{k})\mathrm{d}\boldsymbol{k}.$ (B.2)
Because of Parseval’s relation of the WVD, it follows that
$R_{\mathrm{W}_{f}}(\boldsymbol{0})=\int_{\mathbb{R}^{2N}}\left|\mathrm{W}_{f}(\boldsymbol{z})\right|^{2}\mathrm{d}\boldsymbol{z}=\left\|f\right\|_{2}^{4}.$ (B.3)
Thanks to the conventional convolution and correlation theorems, substituting Eq. (10) yields
$\int_{\mathbb{R}^{2N}}R_{\mathrm{W}_{g},\mathrm{W}_{f}}(-\boldsymbol{k})H_{\mathrm{opt}}(\boldsymbol{k})\mathrm{d}\boldsymbol{k}$
$=$ $\int_{\mathbb{R}^{2N}}\mathcal{F}\left[H_{\mathrm{opt}}\right](\boldsymbol{u})\mathcal{F}\left[\mathrm{W}_{g}\right](\boldsymbol{u})\overline{\mathcal{F}\left[\mathrm{W}_{f}\right](\boldsymbol{u})}\mathrm{d}\boldsymbol{u}$
$=$ $\int_{\mathbb{R}^{2N}}\frac{\varepsilon_{\mathrm{W}_{f},\mathrm{W}_{g}}(\boldsymbol{u})\varepsilon_{\mathrm{W}_{g},\mathrm{W}_{f}}(\boldsymbol{u})}{\varepsilon_{\mathrm{W}_{g}}(\boldsymbol{u})}\mathrm{d}\boldsymbol{u}.$ (B.4)
With Eqs. (B.2), (B.3) and (V-B), we have
$\mathop{\min}\limits_{H(\boldsymbol{z})}\sigma_{\mathrm{MSE}}^{2}=\left\|f\right\|_{2}^{4}-\int_{\mathbb{R}^{2N}}\frac{\varepsilon_{\mathrm{W}_{f},\mathrm{W}_{g}}(\boldsymbol{u})\varepsilon_{\mathrm{W}_{g},\mathrm{W}_{f}}(\boldsymbol{u})}{\varepsilon_{\mathrm{W}_{g}}(\boldsymbol{u})}\mathrm{d}\boldsymbol{u}.$ (B.5)
[…]

### v2

@@
-Adaptive Cohen’s Class Time-Frequency Distribution
+2]Key Laboratory of System Control and Information Processing, Ministry of Education, Shanghai 200240, China
 Abstract
-The fixed kernel function-based Cohen’s class time-frequency distributions (CCTFDs) allow flexibility in denoising for some specific polluted signals.
-Due to the limitation of fixed kernel functions, however, from the view point of filtering they fail to automatically adjust the response according to the change of signal to adapt to different signal characteristics.
-In this letter, we integrate Wiener filter principle and the time-frequency filtering mechanism of CCTFD to design the least-squares adaptive filter method in the Wigner-Ville distribution (WVD) domain, giving birth to the least-squares adaptive filter-based CCTFD whose kernel function can be adjusted with the input signal automatically to achieve the minimum mean-square error denoising in the WVD domain.
+Inspired by the use of adaptive kernel-based Cohen’s class time-frequency distributions (CCTFDs) for cross-term suppression, this paper aims to explore novel adaptive kernel functions for denoising.
+We integrate Wiener filter principle and the time-frequency filtering mechanism of CCTFD to design the least-squares adaptive filter method in the Wigner-Ville distribution (WVD) domain, giving birth to the least-squares adaptive filter-based CCTFD whose kernel function can be adjusted with the input signal automatically to achieve the minimum mean-square error denoising in the WVD domain.
 Some examples are also carried out to demonstrate that the proposed adaptive CCTFD outperforms some state-of-the-arts in noise suppression.
-Index Terms:
+keywords:
 Cohen’s class time-frequency distribution, Convolution, Least-squares adaptive filter, Mean-square error, Power spectral density.
-I Introduction
+1 Introduction
 Cohen’s class time-frequency distribution (CCTFD) [1], also known as the bi-linear kernel function time-frequency distribution, is one of the most representative time-frequency analysis tools of the conventional time-frequency distributions.
@@
 Indeed, it can be regarded as a unified bi-linear time-frequency distribution and has found many applications in seismic exploration, electronic countermeasures, deep-sea detection, spectral imaging, and ultrasonic inspection [9, 10, 11, 12, 13, 14].
-From the view point of signal processing, the integral form of CCTFD can be rewritten as the conventional convolution of the WVD and the kernel function.
-Namely, the CCTFD is a result of smoothing the WVD by using the kernel function.
-The main purpose the smoothing filter is not just only to suppress cross-terms of the WVD, but to achieve non-stationary signal denoising in the WVD domain, and accordingly extract the target signal from background noises or eliminate the impact of noises through the unique reconstruction property of WVD.
-There are a number of attempts to use the filtering instinct of CCTFD to remove noises, giving birth to many useful and effective kernel functions of which the CCTFD works well [15, 16, 17, 18, 19].
-These filtering methods employ fixed kernel functions in dealing with specific input signals, which limits the scope of applications severely.
-To overcome this shortcoming, it seems feasible to explore the adaptive CCTFD whose kernel function can be adjusted with the input signal automatically.
-Adaptive filter is a kind of filter that is able to automatically adjust its parameters or response according to the change of signal to adapt to different signal characteristics [20].
+The integral form of CCTFD can be written as the Fourier transform (FT) of the product of the ambiguity function and the kernel function.
+Cross-terms may arise from the bi-linear nature of CCTFD, causing interference and reducing clarity in signal analysis.
+There are a number of attempts to use the kernel of CCTFD to suppress cross-terms, giving birth to many useful and effective kernel functions of which the CCTFD works well [15, 16, 17, 18, 19].
+However, these methods employ fixed kernel functions in dealing with specific input signals, which limits the scope of applications severely.
+To overcome this shortcoming, Baraniuk et al.
+have proposed a series of methods using adaptive kernel functions to suppress cross-terms [20, 21, 22, 23].
+The adaptive radial Gaussian kernel function-based CCTFD is one of the most famous in this field.
+The integral form of CCTFD can be rewritten as the conventional convolution of the WVD and the kernel function.
+It is important to note that the kernel function mentioned here is a Fourier transform pair with the kernel function described above, as shown in Eq.
+(2).
+From the view point of signal processing, the CCTFD is a result of smoothing the WVD by using the kernel function.
+Inspired by the adaptive kernel function concept proposed by Baraniuk et al., we aim to implement adaptive denoising by designing an adaptive kernel function in the WVD domain.
+This approach will enable us to effectively extract the target signal from background noise and mitigate the impact of noise through the unique reconstruction property of WVD.
+An adaptive filter [24] is a type of filter that automatically adjusts its parameters in response to changes in the signal, making it particularly effective for noise reduction by adapting to varying characteristics of the noise and effectively suppressing it.
 The most celebrated result in this field is Wiener’s result [25] which applies the minimum mean-square error (MSE) criterion to design the least-squares adaptive filter method.
 The Wiener filter has been widely used in addressing practical issues encountered frequently in radar, communications, sonar, biomedicine, and vibration engineering [26, 27, 28, 29, 30, 31, 32].
-The core idea of this letter is to integrate Wiener filter principle and the time-frequency filtering mechanism of CCTFD to investigate the convolution type of CCTFD time-frequency analysis method.
-The main purpose of this letter is to design the least-squares adaptive filter method in the WVD domain, disclose the influence mechanism of the kernel function on the effect of denoising, and establish the adaptive kernel function of CCTFD with the minimum MSE in the WVD domain.
-The main contributions of this letter are summarized as follows:
-  • This letter obtains the least-squares adaptive filter in the WVD domain.
-  • This letter proposes the adaptive CCTFD whose kernel function takes the reversal of the least-squares adaptive filter transfer function in the WVD domain to restore signals from an extreme strong additive noise background.
-  • This letter demonstrates the noise suppression superiority of the proposed adaptive CCTFD over the ordinary Wiener filter and some classical fixed kernel function-based CCTFDs.
-The remainder of this letter is structured as follows.
-In Section II, we recall some necessary background and notation on the convolution type of CCTFD.
-In Section III, we provide the CCTFD-based adaptive filter method for additive noises jamming signals.
-In Section IV, we introduce some examples to validate the effectiveness, reliability and feasibility of the proposed method.
-In Section V, we draw a conclusion.
+The core idea of this paper is to integrate Wiener filter principle and the time-frequency filtering mechanism of CCTFD to investigate the convolution type of CCTFD time-frequency analysis method.
+The main purpose of this paper is to design the least-squares adaptive filter method in the WVD domain, disclose the influence mechanism of the kernel function on the effect of denoising, and establish the adaptive kernel function of CCTFD with the minimum MSE in the WVD domain.
+The main contributions of this paper are summarized as follows:
+  • This paper obtains the least-squares adaptive filter in the WVD domain.
+  • This paper proposes the adaptive CCTFD whose kernel function takes the reversal of the least-squares adaptive filter transfer function in the WVD domain to restore signals from an extreme strong additive noise background.
+  • This paper demonstrates the noise suppression superiority of the proposed adaptive CCTFD over the ordinary Wiener filter, some classical fixed kernel function-based CCTFDs, and the adaptive radial Gaussian kernel function-based CCTFD.
+The remainder of this paper is structured as follows.
+In Section 2, we recall some necessary background and notation on the convolution type of CCTFD.
+In Section 3, we provide the CCTFD-based adaptive filter method for additive noises jamming signals.
+In Section 4, we introduce some examples to validate the effectiveness, reliability and feasibility of the proposed method.
+In Section 5, we draw a conclusion.
 All the technical proofs of our theoretical results are relegated to the appendix parts.
-To facilitate the understanding and clear representation of the various symbols used throughout the paper, we present Table I, which provides a detailed explanation of each symbol and its corresponding meaning.
-TABLE I: Symbol description.
+To facilitate the understanding and clear representation of the various symbols used throughout the paper, we present Table 1, which provides a detailed explanation of each symbol and its corresponding meaning.
+Table 1: Symbol description.
 Symbole Description
@@
 $\mathrm{C}_{g}^{\mathrm{LSAF}}$ the least-squares adaptive filter-based CCTFD for the input signal $g$
-II Convolution type of CCTFD
+2 Convolution type of CCTFD
 The integral form of CCTFD of the function $f(\boldsymbol{x})\in L^{2}(\mathbb{R}^{N})$ reads [1, 33]
-$\mathrm{C}_{f}(\boldsymbol{x},\boldsymbol{w})=$ $\int_{\mathbb{R}^{N}}\int_{\mathbb{R}^{N}}\int_{\mathbb{R}^{N}}f\left(\boldsymbol{y}+\frac{\boldsymbol{\tau}}{2}\right)\overline{f\left(\boldsymbol{y}-\frac{\boldsymbol{\tau}}{2}\right)}$
-×$\times\phi(\boldsymbol{\theta},\boldsymbol{\tau})\mathrm{e}^{-2\pi\mathrm{i}\left(\boldsymbol{\theta}\boldsymbol{x}^{\mathrm{T}}+\boldsymbol{\tau}\boldsymbol{w}^{\mathrm{T}}-\boldsymbol{y}\boldsymbol{\theta}^{\mathrm{T}}\right)}\mathrm{d}\boldsymbol{y}\mathrm{d}\boldsymbol{\tau}\mathrm{d}\boldsymbol{\theta},$ (1)
+$\mathrm{C}_{f}(\boldsymbol{x},\boldsymbol{w})=$ $\int_{\mathbb{R}^{N}}\int_{\mathbb{R}^{N}}\int_{\mathbb{R}^{N}}f\left(\boldsymbol{y}+\frac{\boldsymbol{\tau}}{2}\right)\overline{f\left(\boldsymbol{y}-\frac{\boldsymbol{\tau}}{2}\right)}\phi(\boldsymbol{\theta},\boldsymbol{\tau})\mathrm{e}^{-2\pi\mathrm{i}\left(\boldsymbol{\theta}\boldsymbol{x}^{\mathrm{T}}+\boldsymbol{\tau}\boldsymbol{w}^{\mathrm{T}}-\boldsymbol{y}\boldsymbol{\theta}^{\mathrm{T}}\right)}\mathrm{d}\boldsymbol{y}\mathrm{d}\boldsymbol{\tau}\mathrm{d}\boldsymbol{\theta},$ (1)
 where the superscripts $\mathrm{T}$ and — denote the transpose operator and complex conjugate operator, respectively, and $\phi(\boldsymbol{\theta},\boldsymbol{\tau})$ denotes the kernel function.
-Let $\ast$ and $\mathcal{F}$ be the conventional convolution operator and Fourier operator, respectively. Then, Eq. (II) can be rewritten as
+Let $\ast$ and $\mathcal{F}$ be the conventional convolution operator and Fourier operator, respectively. Then, Eq. (1) can be rewritten as
 $\mathrm{C}_{f}(\boldsymbol{x},\boldsymbol{w})=$ $\left(\mathrm{W}_{f}\ast\Pi\right)(\boldsymbol{x},\boldsymbol{w}),$ (2)
@@
 $=$ $\int_{\mathbb{R}^{N}}\int_{\mathbb{R}^{N}}\phi(\boldsymbol{\theta},\boldsymbol{\tau})\mathrm{e}^{-2\pi\mathrm{i}\left(\boldsymbol{\theta}\boldsymbol{x}^{\mathrm{T}}+\boldsymbol{\tau}\boldsymbol{w}^{\mathrm{T}}\right)}\mathrm{d}\boldsymbol{\theta}\mathrm{d}\boldsymbol{\tau}$ (4)
-denotes the Fourier transform (FT) of the kernel function $\phi(\boldsymbol{\theta},\boldsymbol{\tau})$.
+denotes the FT of the kernel function $\phi(\boldsymbol{\theta},\boldsymbol{\tau})$.
 Eq.
 (2) indicates that the CCTFD is none other than the convolution of the WVD and the (FT version of) kernel function.
-III Least-squares adaptive filter-based CCTFD
+3 Least-squares adaptive filter-based CCTFD
 For a given noise polluted signal $g(\boldsymbol{x})=f(\boldsymbol{x})+n(\boldsymbol{x})$, where $f(\boldsymbol{x})$ and $n(\boldsymbol{x})$ denote the pure signal and the additive noise, respectively, the common tactic of filters is to restore the pure signal as accurately as possible, namely, find the estimate $\widehat{f}(\boldsymbol{x})$ as close as possible to the ideal $f(\boldsymbol{x})$.
@@
 See Appendix B for the detailed derivation of Eq. (12).
-From Eq. (II), it derives that the adaptive optimal kernel function takes the reversal of the least-squares adaptive filter transfer function in the WVD domain, i.e.,
+From Eq. (2), it derives that the adaptive optimal kernel function takes the reversal of the least-squares adaptive filter transfer function in the WVD domain, i.e.,
 $\phi_{\mathrm{opt}}(\boldsymbol{\theta},\boldsymbol{\tau})=\mathcal{F}\left[H_{\mathrm{opt}}\right](-\boldsymbol{\theta},-\boldsymbol{\tau})=\frac{\varepsilon_{\mathrm{W}_{f},\mathrm{W}_{g}}(-\boldsymbol{\theta},-\boldsymbol{\tau})}{\varepsilon_{\mathrm{W}_{g}}(-\boldsymbol{\theta},-\boldsymbol{\tau})}.$ (13)
-Substituting Eq. (13) into Eq. (II) gives the least-squares adaptive filter-based CCTFD
+Substituting Eq. (13) into Eq. (1) gives the least-squares adaptive filter-based CCTFD
 $\mathrm{C}_{g}^{\mathrm{LSAF}}(\boldsymbol{x},\boldsymbol{w})$
-$=$ $\int_{\mathbb{R}^{N}}\int_{\mathbb{R}^{N}}\int_{\mathbb{R}^{N}}g\left(\boldsymbol{y}+\frac{\boldsymbol{\tau}}{2}\right)\overline{g\left(\boldsymbol{y}-\frac{\boldsymbol{\tau}}{2}\right)}$
-×$\times\frac{\varepsilon_{\mathrm{W}_{f},\mathrm{W}_{g}}(-\boldsymbol{\theta},-\boldsymbol{\tau})}{\varepsilon_{\mathrm{W}_{g}}(-\boldsymbol{\theta},-\boldsymbol{\tau})}\mathrm{e}^{-2\pi\mathrm{i}\left(\boldsymbol{\theta}\boldsymbol{x}^{\mathrm{T}}+\boldsymbol{\tau}\boldsymbol{w}^{\mathrm{T}}-\boldsymbol{y}\boldsymbol{\theta}^{\mathrm{T}}\right)}\mathrm{d}\boldsymbol{y}\mathrm{d}\boldsymbol{\tau}\mathrm{d}\boldsymbol{\theta}.$ (14)
+$=$ $\int_{\mathbb{R}^{N}}\int_{\mathbb{R}^{N}}\int_{\mathbb{R}^{N}}g\left(\boldsymbol{y}+\frac{\boldsymbol{\tau}}{2}\right)\overline{g\left(\boldsymbol{y}-\frac{\boldsymbol{\tau}}{2}\right)}\frac{\varepsilon_{\mathrm{W}_{f},\mathrm{W}_{g}}(-\boldsymbol{\theta},-\boldsymbol{\tau})}{\varepsilon_{\mathrm{W}_{g}}(-\boldsymbol{\theta},-\boldsymbol{\tau})}\mathrm{e}^{-2\pi\mathrm{i}\left(\boldsymbol{\theta}\boldsymbol{x}^{\mathrm{T}}+\boldsymbol{\tau}\boldsymbol{w}^{\mathrm{T}}-\boldsymbol{y}\boldsymbol{\theta}^{\mathrm{T}}\right)}\mathrm{d}\boldsymbol{y}\mathrm{d}\boldsymbol{\tau}\mathrm{d}\boldsymbol{\theta}.$ (14)
 It is obvious that the transformation is designed to adapt based on the input signals $f$ and $g$.
@@
 As a result, the adaptive filter dynamically adjusts to varying signal conditions, providing a tailored and efficient filtering process that enhances the clarity and accuracy of the time-frequency representation.
-IV Numerical experiments
+4 Numerical experiments
 This section performs four examples to verify the correctness and effectiveness of the least-squares adaptive filter method in the WVD domain.
-It also compares the denoising effect of the proposed adaptive CCTFD with that of the ordinary Wiener filter and some classical fixed kernel function-based CCTFDs.
-In simulations, the fixed kernel functions are chosen as $\phi(\boldsymbol{\theta},\boldsymbol{\tau})=\cos\left(\frac{\boldsymbol{\theta}\boldsymbol{\tau}^{\mathrm{T}}}{2}\right)$, $\phi(\boldsymbol{\theta},\boldsymbol{\tau})=\mathrm{e}^{\mathrm{i}\frac{\boldsymbol{\theta}\boldsymbol{\tau}^{\mathrm{T}}}{2}}$, $\phi(\boldsymbol{\theta},\boldsymbol{\tau})=\frac{\sin\left(\frac{\boldsymbol{\theta}\boldsymbol{\tau}^{\mathrm{T}}}{2}\right)}{\frac{\boldsymbol{\theta}\boldsymbol{\tau}^{\mathrm{T}}}{2}}$ and $\phi(\boldsymbol{\theta},\boldsymbol{\tau})=\mathrm{e}^{\mathrm{i}\boldsymbol{\theta}\left\|\boldsymbol{\tau}\right\|_{1}}$, (here $\left\|\cdot\right\|_{1}$ denotes the $1$-norm for vectors), corresponding to the Margenau-Hill distribution, the Kirkwood-Rihaczek distribution, the Born-Jordan distribution and the Page distribution, respectively.
+It also compares the denoising effect of the proposed adaptive CCTFD with that of the ordinary Wiener filter, some classical fixed kernel function-based CCTFDs and the adaptive radial Gaussian kernel function-based CCTFD.
+In simulations, the fixed kernel functions are chosen as $\phi(\boldsymbol{\theta},\boldsymbol{\tau})=\cos\left(\frac{\boldsymbol{\theta}\boldsymbol{\tau}^{\mathrm{T}}}{2}\right)$, $\phi(\boldsymbol{\theta},\boldsymbol{\tau})=\mathrm{e}^{\mathrm{i}\frac{\boldsymbol{\theta}\boldsymbol{\tau}^{\mathrm{T}}}{2}}$, $\phi(\boldsymbol{\theta},\boldsymbol{\tau})=\frac{\sin\left(\frac{\boldsymbol{\theta}\boldsymbol{\tau}^{\mathrm{T}}}{2}\right)}{\frac{\boldsymbol{\theta}\boldsymbol{\tau}^{\mathrm{T}}}{2}}$ and $\phi(\boldsymbol{\theta},\boldsymbol{\tau})=\mathrm{e}^{\mathrm{i}\boldsymbol{\theta}\left\|\boldsymbol{\tau}\right\|_{1}}$, (here $\left\|\cdot\right\|_{1}$ denotes the 111-norm for vectors), corresponding to the Margenau-Hill distribution, the Kirkwood-Rihaczek distribution, the Born-Jordan distribution and the Page distribution, respectively.
 For simplicity, we call the filtering methods using fixed kernel functions $\cos\left(\frac{\boldsymbol{\theta}\boldsymbol{\tau}^{\mathrm{T}}}{2}\right)$, $\mathrm{e}^{\mathrm{i}\frac{\boldsymbol{\theta}\boldsymbol{\tau}^{\mathrm{T}}}{2}}$, $\frac{\sin\left(\frac{\boldsymbol{\theta}\boldsymbol{\tau}^{\mathrm{T}}}{2}\right)}{\frac{\boldsymbol{\theta}\boldsymbol{\tau}^{\mathrm{T}}}{2}}$ and $\mathrm{e}^{\mathrm{i}\boldsymbol{\theta}\left\|\boldsymbol{\tau}\right\|_{1}}$ as Margenau-Hill, Kirkwood-Rihaczek, Born-Jordan and Page, respectively.
+Additionally, the filtering method utilizing the adaptive radial Gaussian kernel function-based CCTFD is referred to as the adaptive radial Gaussian kernel.
 Example 1 (Linear frequency-modulated (LFM) signal): The polluted signal is selected for the LFM signal $\mathrm{e}^{2\pi\mathrm{i}\left(\boldsymbol{x}+\frac{\boldsymbol{x}^{2}}{2}\right)}$.
@@
 The sampling frequencies are set as follows: $80\mathrm{Hz}$ for example 1, $100\mathrm{Hz}$ for example 2, $150\mathrm{Hz}$ for example 3 and $175\mathrm{Hz}$ for example 4.
-Figs.
-1(a) and (b) plot respectively the SNR-MSE (logarithm base $10$) and SNR-peak SNR (PSNR) (average of real and imaginary parts) line charts of the estimated LFM siganl using six filtering methods including Margenau-Hill, Kirkwood-Rihaczek, Born-Jordan, Page, Wiener filter and the proposed adaptive CCTFD.
-Figs.
-2(a) and (b) plot respectively the SNR-MSE and SNR-PSNR line charts of the estimate GELFM signals using these six filtering methods.
-Figs.
-3(a) and (b) plot respectively the SNR-MSE and SNR-PSNR line charts of the estimate QFM signals using these six filtering methods.
-Figs.
-4(a) and (b) plot respectively the SNR-MSE and SNR-PSNR line charts of the estimate SFM signals using these six filtering methods.
-It can be seen that the proposed adaptive CCTFD achieves better noise suppression performance than Margenau-Hill, Kirkwood-Rihaczek, Born-Jordan, Page, and Wiener filters under different SNR levels.
-Figure 1: Performance evaluation of six denoising methods for the LFM signal at SNR levels from $-5\mathrm{dB}$ to $5\mathrm{dB}$ under white Gaussian noise: (a) MSE performance of the estimated LFM signals; (b) PSNR performance of the estimated LFM signals.
-Figure 2: Performance evaluation of six denoising methods for the GELFM signal at SNR levels from $-5\mathrm{dB}$ to $5\mathrm{dB}$ under white Gaussian noise: (a) MSE performance of the estimated GELFM signals; (b) PSNR performance of the estimated GELFM signals.
-Figure 3: Performance evaluation of six denoising methods for the QFM signal at SNR levels from $-5\mathrm{dB}$ to $5\mathrm{dB}$ under white Gaussian noise: (a) MSE performance of the estimated QFM signals; (b) PSNR performance of the estimated QFM signals.
-Figure 4: Performance evaluation of six denoising methods for the SFM signal at SNR levels from $-5\mathrm{dB}$ to $5\mathrm{dB}$ under white Gaussian noise: (a) MSE performance of the estimated SFM signals; (b) PSNR performance of the estimated SFM signals.
-Additionally, we investigated the denoising performance of the six methods under conditions of different colored noise.
+Figures 1(a) and (b) plot respectively the SNR-MSE (logarithm base 101010) and SNR-peak SNR (PSNR) (average of real and imaginary parts) line charts of the estimated LFM siganl using seven filtering methods including Margenau-Hill, Kirkwood-Rihaczek, Born-Jordan, Page, adaptive radial Gaussian kernel, Wiener filter and the proposed adaptive CCTFD.
+Figures 2(a) and (b) plot respectively the SNR-MSE and SNR-PSNR line charts of the estimate GELFM signals using these seven filtering methods.
+Figures 3(a) and (b) plot respectively the SNR-MSE and SNR-PSNR line charts of the estimate QFM signals using these seven filtering methods.
+Figures 4(a) and (b) plot respectively the SNR-MSE and SNR-PSNR line charts of the estimate SFM signals using these seven filtering methods.
+It can be seen that the proposed adaptive CCTFD achieves better noise suppression performance than Margenau-Hill, Kirkwood-Rihaczek, Born-Jordan, Page, adaptive radial Gaussian kernel, and Wiener filters under different SNR levels.
+Figure 1: Performance evaluation of seven denoising methods for the LFM signal at SNR levels from $-5\mathrm{dB}$ to $5\mathrm{dB}$ under white Gaussian noise: (a) MSE performance of the estimated LFM signals; (b) PSNR performance of the estimated LFM signals.
+Figure 2: Performance evaluation of seven denoising methods for the GELFM signal at SNR levels from $-5\mathrm{dB}$ to $5\mathrm{dB}$ under white Gaussian noise: (a) MSE performance of the estimated GELFM signals; (b) PSNR performance of the estimated GELFM signals.
+Figure 3: Performance evaluation of seven denoising methods for the QFM signal at SNR levels from $-5\mathrm{dB}$ to $5\mathrm{dB}$ under white Gaussian noise: (a) MSE performance of the estimated QFM signals; (b) PSNR performance of the estimated QFM signals.
+Figure 4: Performance evaluation of seven denoising methods for the SFM signal at SNR levels from $-5\mathrm{dB}$ to $5\mathrm{dB}$ under white Gaussian noise: (a) MSE performance of the estimated SFM signals; (b) PSNR performance of the estimated SFM signals.
+Additionally, we investigated the denoising performance of the seven methods under conditions of different colored noise.
 The four signals were subjected to various types of colored noise: pink noise, blue noise, and red noise.
 The observation interval is set to $[-5\mathrm{s},5\mathrm{s}]$, and the sampling frequencies are set as follows: $30\mathrm{Hz}$ for example 1, $50\mathrm{Hz}$ for example 2, $150\mathrm{Hz}$ for examples 3 and 4.
-Table II presents the MSE and PSNR values of the output signals through the six filters under the influence of three types of colored noise.
+Table 2 presents the MSE and PSNR values of the output signals through the seven filters under the influence of three types of colored noise.
 It can be seen that for blue noise, our proposed adaptive CCTFD significantly outperforms the other methods.
@@
 For pink and red noise, the Wiener filter performs better than ours for QFM and SFM signals.
-However, overall, our method consistently outperforms the fixed kernel function filtering methods.
-TABLE II: The MSE and PSNR of six filtering methods under different colored noise conditions.
+However, overall, our method consistently outperforms the fixed kernel function filtering methods and the adaptive radial Gaussian kernel function filtering method.
+Table 2: The MSE and PSNR of seven filtering methods under different colored noise conditions.
 Example LFM GELFM QFM SFM
 MSE (Mean-Square Error)
-Pink noise Margenau-Hill $-0.0016$ $-0.1592$ $-0.0513$ $-0.1319$
-Kirkwood-Rihaczek $0.0676$ $-0.0050$ $-0.0330$ $-0.1200$
-Born-Jordan $-0.1582$ $-0.3053$ $-0.1123$ $-0.2134$
-Page $0.0996$ $-0.0360$ $-0.0201$ $-0.0582$
-Wiener filter $-0.4985$ $-0.6911$ $\mathbf{-0.4699}$ $\mathbf{-0.4268}$
-Adaptive CCTFD $\mathbf{-1.5318}$ $\mathbf{-0.9807}$ $-0.3169$ $-0.3369$
-Blue noise Margenau-Hill $0.1437$ $-0.4563$ $-0.0637$ $-0.3730$
-Kirkwood-Rihaczek $0.3580$ $-0.3711$ $0.0063$ $-0.2404$
-Born-Jordan $-0.1057$ $-0.5265$ $-0.2016$ $-0.9358$
-Page $0.3281$ $-0.3355$ $0.0239$ $0.0682$
-Wiener filter $-0.4676$ $-0.6632$ $-0.5208$ $-1.0831$
+Pink noise Margenau-Hill 0.0016-0.0016- 0.0016 0.1592-0.1592- 0.1592 0.0513-0.0513- 0.0513 0.1319-0.1319- 0.1319
+Kirkwood-Rihaczek 0.06760.06760.0676 0.0050-0.0050- 0.0050 0.0330-0.0330- 0.0330 0.1200-0.1200- 0.1200
+Born-Jordan 0.1582-0.1582- 0.1582 0.3053-0.3053- 0.3053 0.1123-0.1123- 0.1123 0.2134-0.2134- 0.2134
+Page 0.09960.09960.0996 0.0360-0.0360- 0.0360 0.0201-0.0201- 0.0201 0.0582-0.0582- 0.0582
+Adaptive radial Gaussian kernel 0.0451-0.0451- 0.0451 0.4237-0.4237- 0.4237 0.0662-0.0662- 0.0662 0.0510-0.0510- 0.0510
+Wiener filter 0.4985-0.4985- 0.4985 0.6911-0.6911- 0.6911 $\mathbf{-0.4699}$ $\mathbf{-0.4268}$
+Adaptive CCTFD $\mathbf{-1.5318}$ $\mathbf{-0.9807}$ 0.3169-0.3169- 0.3169 0.3369-0.3369- 0.3369
+Blue noise Margenau-Hill 0.14370.14370.1437 0.4563-0.4563- 0.4563 0.0637-0.0637- 0.0637 0.3730-0.3730- 0.3730
+Kirkwood-Rihaczek 0.35800.35800.3580 0.3711-0.3711- 0.3711 0.00630.00630.0063 0.2404-0.2404- 0.2404
+Born-Jordan 0.1057-0.1057- 0.1057 0.5265-0.5265- 0.5265 0.2016-0.2016- 0.2016 0.9358-0.9358- 0.9358
+Page 0.32810.32810.3281 0.3355-0.3355- 0.3355 0.02390.02390.0239 0.06820.06820.0682
+Adaptive radial Gaussian kernel 0.0481-0.0481- 0.0481 0.4892-0.4892- 0.4892 0.0671-0.0671- 0.0671 0.0191-0.0191- 0.0191
+Wiener filter 0.4676-0.4676- 0.4676 0.6632-0.6632- 0.6632 0.5208-0.5208- 0.5208 1.0831-1.0831- 1.0831
 Adaptive CCTFD $\mathbf{-1.2999}$ $\mathbf{-0.9581}$ $\mathbf{-0.7917}$ $\mathbf{-1.6128}$
-Red noise Margenau-Hill $0.0655$ $-0.3004$ $-0.0624$ $-0.2233$
-Kirkwood-Rihaczek $0.1243$ $-0.2567$ $-0.0475$ $-0.2145$
-Born-Jordan $-0.0907$ $-0.3810$ $-0.1369$ $-0.3623$
-Page $0.1720$ $-0.2790$ $-0.0082$ $-0.0851$
-Wiener filter $-0.8039$ $-1.2798$ $\mathbf{-0.9489}$ $\mathbf{-0.7804}$
-Adaptive CCTFD $\mathbf{-1.7448}$ $\mathbf{-1.3028}$ $-0.4032$ $-0.3919$
+Red noise Margenau-Hill 0.06550.06550.0655 0.3004-0.3004- 0.3004 0.0624-0.0624- 0.0624 0.2233-0.2233- 0.2233
+Kirkwood-Rihaczek 0.12430.12430.1243 0.2567-0.2567- 0.2567 0.0475-0.0475- 0.0475 0.2145-0.2145- 0.2145
+Born-Jordan 0.0907-0.0907- 0.0907 0.3810-0.3810- 0.3810 0.1369-0.1369- 0.1369 0.3623-0.3623- 0.3623
+Page 0.17200.17200.1720 0.2790-0.2790- 0.2790 0.0082-0.0082- 0.0082 0.0851-0.0851- 0.0851
+Adaptive radial Gaussian kernel 0.01050.01050.0105 0.3455-0.3455- 0.3455 0.0743-0.0743- 0.0743 0.0395-0.0395- 0.0395
+Wiener filter 0.8039-0.8039- 0.8039 1.2798-1.2798- 1.2798 $\mathbf{-0.9489}$ $\mathbf{-0.7804}$
+Adaptive CCTFD $\mathbf{-1.7448}$ $\mathbf{-1.3028}$ 0.4032-0.4032- 0.4032 0.3919-0.3919- 0.3919
 PSNR (Peak Signal-to-Noise Ratio)
-Pink noise Margenau-Hill $3.0323$ $5.2766$ $3.5237$ $4.3306$
-Kirkwood-Rihaczek $2.3400$ $3.7410$ $3.3431$ $4.2110$
-Born-Jordan $4.6190$ $7.0716$ $4.1339$ $5.1494$
-Page $2.0433$ $3.7370$ $3.2109$ $3.5963$
-Wiener filter $8.8314$ $10.3192$ $\mathbf{8.0556}$ $\mathbf{7.8716}$
-Adaptive CCTFD $\mathbf{18.5390}$ $\mathbf{12.8503}$ $6.1978$ $6.5799$
-Blue noise Margenau-Hill $1.5865$ $7.6228$ $3.6468$ $7.2377$
-Kirkwood-Rihaczek $-0.5454$ $6.8108$ $2.9469$ $5.4278$
-Born-Jordan $4.1557$ $8.3757$ $5.0342$ $13.0179$
-Page $-0.2493$ $6.4764$ $2.7708$ $2.3297$
-Wiener filter $8.1082$ $9.8937$ $8.6899$ $14.4000$
+Pink noise Margenau-Hill 3.03233.03233.0323 5.27665.27665.2766 3.52373.52373.5237 4.33064.33064.3306
+Kirkwood-Rihaczek 2.34002.34002.3400 3.74103.74103.7410 3.34313.34313.3431 4.21104.21104.2110
+Born-Jordan 4.61904.61904.6190 7.07167.07167.0716 4.13394.13394.1339 5.14945.14945.1494
+Page 2.04332.04332.0433 3.73703.73703.7370 3.21093.21093.2109 3.59633.59633.5963
+Adaptive radial Gaussian kernel 3.47583.47583.4758 7.31207.31207.3120 3.67363.67363.6736 3.52163.52163.5216
+Wiener filter 8.83148.83148.8314 10.319210.319210.3192 $\mathbf{8.0556}$ $\mathbf{7.8716}$
+Adaptive CCTFD $\mathbf{18.5390}$ $\mathbf{12.8503}$ 6.19786.19786.1978 6.57996.57996.5799
+Blue noise Margenau-Hill 1.58651.58651.5865 7.62287.62287.6228 3.64683.64683.6468 7.23777.23777.2377
+Kirkwood-Rihaczek 0.5454-0.5454- 0.5454 6.81086.81086.8108 2.94692.94692.9469 5.42785.42785.4278
+Born-Jordan 4.15574.15574.1557 8.37578.37578.3757 5.03425.03425.0342 13.017913.017913.0179
+Page 0.2493-0.2493- 0.2493 6.47646.47646.4764 2.77082.77082.7708 2.32972.32972.3297
+Adaptive radial Gaussian kernel 3.50463.50463.5046 7.90757.90757.9075 3.68133.68133.6813 3.20163.20163.2016
+Wiener filter 8.10828.10828.1082 9.89379.89379.8937 8.68998.68998.6899 14.400014.400014.4000
 Adaptive CCTFD $\mathbf{16.0114}$ $\mathbf{12.6000}$ $\mathbf{10.9354}$ $\mathbf{19.6753}$
-Red noise Margenau-Hill $2.3888$ $6.3998$ $3.6360$ $5.2562$
-Kirkwood-Rihaczek $1.7774$ $5.9456$ $3.4881$ $5.1559$
-Born-Jordan $4.0111$ $7.4167$ $4.3805$ $6.6348$
-Page $1.3606$ $5.9863$ $3.0925$ $3.8628$
-Wiener filter $11.8832$ $\mathbf{16.7052}$ $\mathbf{13.1156}$ $\mathbf{11.9071}$
-Adaptive CCTFD $\mathbf{20.4654}$ $16.1495$ $7.0496$ $7.2056$
-V Conclusion
+Red noise Margenau-Hill 2.38882.38882.3888 6.39986.39986.3998 3.63603.63603.6360 5.25625.25625.2562
+Kirkwood-Rihaczek 1.77741.77741.7774 5.94565.94565.9456 3.48813.48813.4881 5.15595.15595.1559
+Born-Jordan 4.01114.01114.0111 7.41677.41677.4167 4.38054.38054.3805 6.63486.63486.6348
+Page 1.36061.36061.3606 5.98635.98635.9863 3.09253.09253.0925 3.86283.86283.8628
+Adaptive radial Gaussian kernel 2.99442.99442.9944 6.62676.62676.6267 3.75523.75523.7552 3.40613.40613.4061
+Wiener filter 11.883211.883211.8832 $\mathbf{16.7052}$ $\mathbf{13.1156}$ $\mathbf{11.9071}$
+Adaptive CCTFD $\mathbf{20.4654}$ 16.149516.149516.1495 7.04967.04967.0496 7.20567.20567.2056
+5 Conclusion
 The convolution type of CCTFD time-frequency analysis method suitable for the adaptive filter denoising for additive noises jamming signals under the condition of low SNR was established, giving rise to the adaptive kernel function which minimizes the MSE in the WVD domain.
 The proposed adaptive CCTFD can automatically adjust its kernel function according to the change of signal to adapt to different signal characteristics.
-It turns out that its denoising effect is better than that of some classical fixed kernel function-based CCTFDs, including the Margenau-Hill distribution, the Kirkwood-Rihaczek distribution, the Born-Jordan distribution, and the Page distribution.
+It turns out that its denoising effect is superior not only to some classical fixed kernel function-based CCTFDs, including the Margenau-Hill distribution, the Kirkwood-Rihaczek distribution, the Born-Jordan distribution, and the Page distribution, but also to the adaptive radial Gaussian kernel function-based CCTFD.
 Under white Gaussian noise, our method outperforms the ordinary Wiener filter, while under colored noise, it performs comparably to the Wiener filter.
-Appendix
-In this section, we prove all the theoretical results in the letter.
-V-A Proof of Eq. (10)
+[…]
+Acknowledgements This work was supported by the National Natural Science Foundation of China (Grant No 61901223), the Jiangsu Planned Projects for Postdoctoral Research Funds (Grant No 2021K205B), the Foundation of Key Laboratory of System Control and Information Processing, Ministry of Education (Grant No Scip20240121), and the Foundation of Key Laboratory of Computational Science and Application of Hainan Province (Grant No JSKX202401).
+Appendix A Proof of Eq. (10)
 By using the orthogonal principle
@@
 […]
-V-B Proof of Eq. (12)
+Appendix B Proof of Eq. (12)
 The minimize MSE takes
@@
 $=$ $\int_{\mathbb{R}^{2N}}\frac{\varepsilon_{\mathrm{W}_{f},\mathrm{W}_{g}}(\boldsymbol{u})\varepsilon_{\mathrm{W}_{g},\mathrm{W}_{f}}(\boldsymbol{u})}{\varepsilon_{\mathrm{W}_{g}}(\boldsymbol{u})}\mathrm{d}\boldsymbol{u}.$ (B.4)
-With Eqs. (B.2), (B.3) and (V-B), we have
+With Eqs. (B.2), (B.3) and (B), we have
 $\mathop{\min}\limits_{H(\boldsymbol{z})}\sigma_{\mathrm{MSE}}^{2}=\left\|f\right\|_{2}^{4}-\int_{\mathbb{R}^{2N}}\frac{\varepsilon_{\mathrm{W}_{f},\mathrm{W}_{g}}(\boldsymbol{u})\varepsilon_{\mathrm{W}_{g},\mathrm{W}_{f}}(\boldsymbol{u})}{\varepsilon_{\mathrm{W}_{g}}(\boldsymbol{u})}\mathrm{d}\boldsymbol{u}.$ (B.5)

### v3

@@
 We integrate Wiener filter principle and the time-frequency filtering mechanism of CCTFD to design the least-squares adaptive filter method in the Wigner-Ville distribution (WVD) domain, giving birth to the least-squares adaptive filter-based CCTFD whose kernel function can be adjusted with the input signal automatically to achieve the minimum mean-square error denoising in the WVD domain.
-Some examples are also carried out to demonstrate that the proposed adaptive CCTFD outperforms some state-of-the-arts in noise suppression.
+Numerical experiments on both simulated and real-world signals demonstrate that the proposed adaptive CCTFD outperforms some state-of-the-art methods in noise suppression.
 keywords:
@@
 1 Introduction
-Cohen’s class time-frequency distribution (CCTFD) [1], also known as the bi-linear kernel function time-frequency distribution, is one of the most representative time-frequency analysis tools of the conventional time-frequency distributions.
-It includes particular cases the Wigner-Ville distribution (WVD) [2], the Choi-Williams distribution [3], the Kirkwood-Rihaczek distribution [4], the Born-Jordan distribution [5], the Zhao-Atlas-Marks distribution [6], the Margenau-Hill distribution [7], and the Page distribution [8].
-Indeed, it can be regarded as a unified bi-linear time-frequency distribution and has found many applications in seismic exploration, electronic countermeasures, deep-sea detection, spectral imaging, and ultrasonic inspection [9, 10, 11, 12, 13, 14].
+Cohen’s class time-frequency distribution (CCTFD) [Coh95], also known as the bi-linear kernel function time-frequency distribution, is one of the most representative time-frequency analysis tools of the conventional time-frequency distributions.
+It includes particular cases the Wigner-Ville distribution (WVD) [wigner1932quantum], the Choi-Williams distribution [zhao2021research], the Kirkwood-Rihaczek distribution [rihaczek1968signal], the Born-Jordan distribution [quan2021novel], the Zhao-Atlas-Marks distribution [zhao1990use], the Margenau-Hill distribution [liang2024spatial], and the Page distribution [xia1996generalized].
+Indeed, it can be regarded as a unified bi-linear time-frequency distribution and has found many applications in seismic exploration, electronic countermeasures, deep-sea detection, spectral imaging, and ultrasonic inspection [brynolfsson2021time, wang2021matching, xu2023dtftcnet, vitor2023induction, han2021passive, ding2023regional].
+In addition to traditional CCTFD and their variants, deep learning-based adaptive time-frequency analysis methods have attracted attention for their ability to learn optimal representations from data.
+These methods are useful for signal filtering, denoising, and separation, and have shown effectiveness in tasks such as video, speech, and radar signal analysis [guo2023deep, pan2022tfa, li2023incipient, samavat2022deep, hamdaoui2023improved, zhang2024adaptive].
 The integral form of CCTFD can be written as the Fourier transform (FT) of the product of the ambiguity function and the kernel function.
 Cross-terms may arise from the bi-linear nature of CCTFD, causing interference and reducing clarity in signal analysis.
-There are a number of attempts to use the kernel of CCTFD to suppress cross-terms, giving birth to many useful and effective kernel functions of which the CCTFD works well [15, 16, 17, 18, 19].
+There are a number of attempts to use the kernel of CCTFD to suppress cross-terms, giving birth to many useful and effective kernel functions of which the CCTFD works well [qu2020radar, yao2012reducing, lopac2021detection, jadhav2022automated, li2023time].
 However, these methods employ fixed kernel functions in dealing with specific input signals, which limits the scope of applications severely.
 To overcome this shortcoming, Baraniuk et al.
-have proposed a series of methods using adaptive kernel functions to suppress cross-terms [20, 21, 22, 23].
+have proposed a series of methods using adaptive kernel functions to suppress cross-terms [baraniuk1991radially, 212733, 258128, jones1995adaptive].
 The adaptive radial Gaussian kernel function-based CCTFD is one of the most famous in this field.
+However, existing adaptive kernel methods have been primarily developed with the objective of cross-term suppression rather than signal denoising.
+These approaches adapt the kernel shape based on the local structure of the ambiguity function to enhance time-frequency concentration and reduce interference between components.
+While such techniques improve the readability of time-frequency representations, they do not explicitly address the problem of noise suppression or signal reconstruction optimality.
+Consequently, when the signal is heavily contaminated by noise, conventional adaptive kernel methods often fail to achieve the desired filtering performance.
+To overcome this limitation, we reformulate the adaptive kernel design problem from the perspective of optimal time-frequency filtering.
 The integral form of CCTFD can be rewritten as the conventional convolution of the WVD and the kernel function.
@@
 This approach will enable us to effectively extract the target signal from background noise and mitigate the impact of noise through the unique reconstruction property of WVD.
-An adaptive filter [24] is a type of filter that automatically adjusts its parameters in response to changes in the signal, making it particularly effective for noise reduction by adapting to varying characteristics of the noise and effectively suppressing it.
-The most celebrated result in this field is Wiener’s result [25] which applies the minimum mean-square error (MSE) criterion to design the least-squares adaptive filter method.
-The Wiener filter has been widely used in addressing practical issues encountered frequently in radar, communications, sonar, biomedicine, and vibration engineering [26, 27, 28, 29, 30, 31, 32].
+An adaptive filter [diniz1997adaptive] is a type of filter that automatically adjusts its parameters in response to changes in the signal, making it particularly effective for noise reduction by adapting to varying characteristics of the noise and effectively suppressing it.
+The most celebrated result in this field is Wiener’s result [wiener1949extrapolation] which applies the minimum mean-square error (MSE) criterion to design the least-squares adaptive filter method.
+The Wiener filter has been widely used in addressing practical issues encountered frequently in radar, communications, sonar, biomedicine, and vibration engineering [ceccato2020spatial, baudais2023doppler, plabst2020wiener, song2020two, chang2022tensor, qi2021wiener, vered2024parallel].
+Unlike previous adaptive kernel designs focused on cross-term suppression, the proposed method establishes a framework that bridges time-frequency representation and adaptive filtering theory, thereby providing a fundamental advance in both interpretability and performance.
 The core idea of this paper is to integrate Wiener filter principle and the time-frequency filtering mechanism of CCTFD to investigate the convolution type of CCTFD time-frequency analysis method.
@@
   • This paper obtains the least-squares adaptive filter in the WVD domain.
-  • This paper proposes the adaptive CCTFD whose kernel function takes the reversal of the least-squares adaptive filter transfer function in the WVD domain to restore signals from an extreme strong additive noise background.
+  • This paper proposes the adaptive CCTFD whose kernel function takes the reversal of the least-squares adaptive filter transfer function in the WVD domain to restore signals from an extremely strong additive noise background.
   • This paper demonstrates the noise suppression superiority of the proposed adaptive CCTFD over the ordinary Wiener filter, some classical fixed kernel function-based CCTFDs, and the adaptive radial Gaussian kernel function-based CCTFD.
@@
 In Section 3, we provide the CCTFD-based adaptive filter method for additive noises jamming signals.
-In Section 4, we introduce some examples to validate the effectiveness, reliability and feasibility of the proposed method.
+In Section 4, we introduce numerical examples and real-world data experiments to validate the effectiveness, reliability, and feasibility of the proposed method.
 In Section 5, we draw a conclusion.
@@
 $\varepsilon$ PSD
-$\mathrm{C}_{g}^{\mathrm{LSAF}}$ the least-squares adaptive filter-based CCTFD for the input signal $g$
+$\mathrm{C}_{g}^{\mathrm{LSAF}}$ the least-squares adaptive filter-based CCTFD
+for the input signal $g$
 2 Convolution type of CCTFD
-The integral form of CCTFD of the function $f(\boldsymbol{x})\in L^{2}(\mathbb{R}^{N})$ reads [1, 33]
-$\mathrm{C}_{f}(\boldsymbol{x},\boldsymbol{w})=$ $\int_{\mathbb{R}^{N}}\int_{\mathbb{R}^{N}}\int_{\mathbb{R}^{N}}f\left(\boldsymbol{y}+\frac{\boldsymbol{\tau}}{2}\right)\overline{f\left(\boldsymbol{y}-\frac{\boldsymbol{\tau}}{2}\right)}\phi(\boldsymbol{\theta},\boldsymbol{\tau})\mathrm{e}^{-2\pi\mathrm{i}\left(\boldsymbol{\theta}\boldsymbol{x}^{\mathrm{T}}+\boldsymbol{\tau}\boldsymbol{w}^{\mathrm{T}}-\boldsymbol{y}\boldsymbol{\theta}^{\mathrm{T}}\right)}\mathrm{d}\boldsymbol{y}\mathrm{d}\boldsymbol{\tau}\mathrm{d}\boldsymbol{\theta},$ (1)
+The integral form of CCTFD of the function $f(\boldsymbol{x})\in L^{2}(\mathbb{R}^{N})$ reads [Coh95, ozaktas1996effect]
+$\mathrm{C}_{f}(\boldsymbol{x},\boldsymbol{w})=$ $\int_{\mathbb{R}^{N}}\int_{\mathbb{R}^{N}}\int_{\mathbb{R}^{N}}f\left(\boldsymbol{y}+\frac{\boldsymbol{\tau}}{2}\right)\overline{f\left(\boldsymbol{y}-\frac{\boldsymbol{\tau}}{2}\right)}$
+$\times\phi(\boldsymbol{\theta},\boldsymbol{\tau})\mathrm{e}^{-2\pi\mathrm{i}\left(\boldsymbol{\theta}\boldsymbol{x}^{\mathrm{T}}+\boldsymbol{\tau}\boldsymbol{w}^{\mathrm{T}}-\boldsymbol{y}\boldsymbol{\theta}^{\mathrm{T}}\right)}\mathrm{d}\boldsymbol{y}\mathrm{d}\boldsymbol{\tau}\mathrm{d}\boldsymbol{\theta},$ (1)
 where the superscripts $\mathrm{T}$ and — denote the transpose operator and complex conjugate operator, respectively, and $\phi(\boldsymbol{\theta},\boldsymbol{\tau})$ denotes the kernel function.
@@
 $H_{\mathrm{opt}}(\boldsymbol{z})=\mathop{\arg\min}\limits_{H(\boldsymbol{z})}\sigma_{\mathrm{MSE}}^{2}.$ (9)
-By using the orthogonal principle [34], the stationary assumption and the conventional convolution and correlation theorems to establish, simplify and solve the Wiener-Hopf equation, respectively, the least-squares adaptive filter transfer function in the WVD domain reads
+By using the orthogonal principle [proakis1985probability], the stationary assumption and the conventional convolution and correlation theorems to establish, simplify and solve the Wiener-Hopf equation, respectively, the least-squares adaptive filter transfer function in the WVD domain reads
 $\mathcal{F}\left[H_{\mathrm{opt}}\right](\boldsymbol{u})=\frac{\varepsilon_{\mathrm{W}_{f},\mathrm{W}_{g}}(\boldsymbol{u})}{\varepsilon_{\mathrm{W}_{g}}(\boldsymbol{u})},$ (10)
@@
 $\mathrm{C}_{g}^{\mathrm{LSAF}}(\boldsymbol{x},\boldsymbol{w})$
-$=$ $\int_{\mathbb{R}^{N}}\int_{\mathbb{R}^{N}}\int_{\mathbb{R}^{N}}g\left(\boldsymbol{y}+\frac{\boldsymbol{\tau}}{2}\right)\overline{g\left(\boldsymbol{y}-\frac{\boldsymbol{\tau}}{2}\right)}\frac{\varepsilon_{\mathrm{W}_{f},\mathrm{W}_{g}}(-\boldsymbol{\theta},-\boldsymbol{\tau})}{\varepsilon_{\mathrm{W}_{g}}(-\boldsymbol{\theta},-\boldsymbol{\tau})}\mathrm{e}^{-2\pi\mathrm{i}\left(\boldsymbol{\theta}\boldsymbol{x}^{\mathrm{T}}+\boldsymbol{\tau}\boldsymbol{w}^{\mathrm{T}}-\boldsymbol{y}\boldsymbol{\theta}^{\mathrm{T}}\right)}\mathrm{d}\boldsymbol{y}\mathrm{d}\boldsymbol{\tau}\mathrm{d}\boldsymbol{\theta}.$ (14)
+$=$ $\int_{\mathbb{R}^{N}}\int_{\mathbb{R}^{N}}\int_{\mathbb{R}^{N}}g\left(\boldsymbol{y}+\frac{\boldsymbol{\tau}}{2}\right)\overline{g\left(\boldsymbol{y}-\frac{\boldsymbol{\tau}}{2}\right)}$
+$\times\frac{\varepsilon_{\mathrm{W}_{f},\mathrm{W}_{g}}(-\boldsymbol{\theta},-\boldsymbol{\tau})}{\varepsilon_{\mathrm{W}_{g}}(-\boldsymbol{\theta},-\boldsymbol{\tau})}\mathrm{e}^{-2\pi\mathrm{i}\left(\boldsymbol{\theta}\boldsymbol{x}^{\mathrm{T}}+\boldsymbol{\tau}\boldsymbol{w}^{\mathrm{T}}-\boldsymbol{y}\boldsymbol{\theta}^{\mathrm{T}}\right)}\mathrm{d}\boldsymbol{y}\mathrm{d}\boldsymbol{\tau}\mathrm{d}\boldsymbol{\theta}.$ (14)
 It is obvious that the transformation is designed to adapt based on the input signals $f$ and $g$.
@@
 This section performs four examples to verify the correctness and effectiveness of the least-squares adaptive filter method in the WVD domain.
-It also compares the denoising effect of the proposed adaptive CCTFD with that of the ordinary Wiener filter, some classical fixed kernel function-based CCTFDs and the adaptive radial Gaussian kernel function-based CCTFD.
-In simulations, the fixed kernel functions are chosen as $\phi(\boldsymbol{\theta},\boldsymbol{\tau})=\cos\left(\frac{\boldsymbol{\theta}\boldsymbol{\tau}^{\mathrm{T}}}{2}\right)$, $\phi(\boldsymbol{\theta},\boldsymbol{\tau})=\mathrm{e}^{\mathrm{i}\frac{\boldsymbol{\theta}\boldsymbol{\tau}^{\mathrm{T}}}{2}}$, $\phi(\boldsymbol{\theta},\boldsymbol{\tau})=\frac{\sin\left(\frac{\boldsymbol{\theta}\boldsymbol{\tau}^{\mathrm{T}}}{2}\right)}{\frac{\boldsymbol{\theta}\boldsymbol{\tau}^{\mathrm{T}}}{2}}$ and $\phi(\boldsymbol{\theta},\boldsymbol{\tau})=\mathrm{e}^{\mathrm{i}\boldsymbol{\theta}\left\|\boldsymbol{\tau}\right\|_{1}}$, (here $\left\|\cdot\right\|_{1}$ denotes the 111-norm for vectors), corresponding to the Margenau-Hill distribution, the Kirkwood-Rihaczek distribution, the Born-Jordan distribution and the Page distribution, respectively.
+The experiments include both simulated signals and real-world data from the Radar Characterisation (RadChar) dataset [Huang23Multi].
+The denoising performance of the proposed adaptive CCTFD is compared with that of the ordinary Wiener filter, some classical fixed kernel function-based CCTFDs and the adaptive radial Gaussian kernel function-based CCTFD.
+In simulations, the fixed kernel functions are chosen as $\phi(\boldsymbol{\theta},\boldsymbol{\tau})=\cos\left(\frac{\boldsymbol{\theta}\boldsymbol{\tau}^{\mathrm{T}}}{2}\right)$, $\phi(\boldsymbol{\theta},\boldsymbol{\tau})=\mathrm{e}^{\mathrm{i}\frac{\boldsymbol{\theta}\boldsymbol{\tau}^{\mathrm{T}}}{2}}$, $\phi(\boldsymbol{\theta},\boldsymbol{\tau})=\frac{\sin\left(\frac{\boldsymbol{\theta}\boldsymbol{\tau}^{\mathrm{T}}}{2}\right)}{\frac{\boldsymbol{\theta}\boldsymbol{\tau}^{\mathrm{T}}}{2}}$ and $\phi(\boldsymbol{\theta},\boldsymbol{\tau})=\mathrm{e}^{\mathrm{i}\boldsymbol{\theta}\left\|\boldsymbol{\tau}\right\|_{1}}$, (here $\left\|\cdot\right\|_{1}$ denotes the $1$-norm for vectors), corresponding to the Margenau-Hill distribution, the Kirkwood-Rihaczek distribution, the Born-Jordan distribution and the Page distribution, respectively.
 For simplicity, we call the filtering methods using fixed kernel functions $\cos\left(\frac{\boldsymbol{\theta}\boldsymbol{\tau}^{\mathrm{T}}}{2}\right)$, $\mathrm{e}^{\mathrm{i}\frac{\boldsymbol{\theta}\boldsymbol{\tau}^{\mathrm{T}}}{2}}$, $\frac{\sin\left(\frac{\boldsymbol{\theta}\boldsymbol{\tau}^{\mathrm{T}}}{2}\right)}{\frac{\boldsymbol{\theta}\boldsymbol{\tau}^{\mathrm{T}}}{2}}$ and $\mathrm{e}^{\mathrm{i}\boldsymbol{\theta}\left\|\boldsymbol{\tau}\right\|_{1}}$ as Margenau-Hill, Kirkwood-Rihaczek, Born-Jordan and Page, respectively.
@@
 Example 4 (Sinusoidal frequency-modulated (SFM) signal): The polluted signal is selected for the SFM signal $\mathrm{e}^{\mathrm{i}\left[1.3\pi\boldsymbol{x}+2\sin(0.6\pi\boldsymbol{x})\right]}$.
-In examples 1–4, additive white Gaussian noise is introduced to the signals, with the SNR ranging from $-5\mathrm{dB}$ to $5\mathrm{dB}$ and the observation interval is set to $[-5\mathrm{s},5\mathrm{s}]$.
+In examples 1–4, additive white Gaussian noise is introduced to the signals, with the SNR ranging from $5\mathrm{dB}$ to $5\mathrm{dB}$ and the observation interval is set to $5\mathrm{s},5\mathrm{s}]$.
 The sampling frequencies are set as follows: $80\mathrm{Hz}$ for example 1, $100\mathrm{Hz}$ for example 2, $150\mathrm{Hz}$ for example 3 and $175\mathrm{Hz}$ for example 4.
-Figures 1(a) and (b) plot respectively the SNR-MSE (logarithm base 101010) and SNR-peak SNR (PSNR) (average of real and imaginary parts) line charts of the estimated LFM siganl using seven filtering methods including Margenau-Hill, Kirkwood-Rihaczek, Born-Jordan, Page, adaptive radial Gaussian kernel, Wiener filter and the proposed adaptive CCTFD.
-Figures 2(a) and (b) plot respectively the SNR-MSE and SNR-PSNR line charts of the estimate GELFM signals using these seven filtering methods.
-Figures 3(a) and (b) plot respectively the SNR-MSE and SNR-PSNR line charts of the estimate QFM signals using these seven filtering methods.
-Figures 4(a) and (b) plot respectively the SNR-MSE and SNR-PSNR line charts of the estimate SFM signals using these seven filtering methods.
+Figs 1(a) and (b) plot respectively the SNR-MSE (logarithm base $10$) and SNR-peak SNR (PSNR) (average of real and imaginary parts) line charts of the estimated LFM signal using seven filtering methods including Margenau-Hill, Kirkwood-Rihaczek, Born-Jordan, Page, adaptive radial Gaussian kernel, Wiener filter and the proposed adaptive CCTFD.
+Figs 2(a) and (b) plot respectively the SNR-MSE and SNR-PSNR line charts of the estimate GELFM signals using these seven filtering methods.
+Figs 3(a) and (b) plot respectively the SNR-MSE and SNR-PSNR line charts of the estimate QFM signals using these seven filtering methods.
+Figs 4(a) and (b) plot respectively the SNR-MSE and SNR-PSNR line charts of the estimate SFM signals using these seven filtering methods.
 It can be seen that the proposed adaptive CCTFD achieves better noise suppression performance than Margenau-Hill, Kirkwood-Rihaczek, Born-Jordan, Page, adaptive radial Gaussian kernel, and Wiener filters under different SNR levels.
-Figure 1: Performance evaluation of seven denoising methods for the LFM signal at SNR levels from $-5\mathrm{dB}$ to $5\mathrm{dB}$ under white Gaussian noise: (a) MSE performance of the estimated LFM signals; (b) PSNR performance of the estimated LFM signals.
-Figure 2: Performance evaluation of seven denoising methods for the GELFM signal at SNR levels from $-5\mathrm{dB}$ to $5\mathrm{dB}$ under white Gaussian noise: (a) MSE performance of the estimated GELFM signals; (b) PSNR performance of the estimated GELFM signals.
-Figure 3: Performance evaluation of seven denoising methods for the QFM signal at SNR levels from $-5\mathrm{dB}$ to $5\mathrm{dB}$ under white Gaussian noise: (a) MSE performance of the estimated QFM signals; (b) PSNR performance of the estimated QFM signals.
-Figure 4: Performance evaluation of seven denoising methods for the SFM signal at SNR levels from $-5\mathrm{dB}$ to $5\mathrm{dB}$ under white Gaussian noise: (a) MSE performance of the estimated SFM signals; (b) PSNR performance of the estimated SFM signals.
-Additionally, we investigated the denoising performance of the seven methods under conditions of different colored noise.
+Figure 1: Performance evaluation of seven denoising methods for the LFM signal at SNR levels from $5\mathrm{dB}$ to $5\mathrm{dB}$ under white Gaussian noise: (a) MSE performance of the estimated LFM signals; (b) PSNR performance of the estimated LFM signals.
+Figure 2: Performance evaluation of seven denoising methods for the GELFM signal at SNR levels from $5\mathrm{dB}$ to $5\mathrm{dB}$ under white Gaussian noise: (a) MSE performance of the estimated GELFM signals; (b) PSNR performance of the estimated GELFM signals.
+Figure 3: Performance evaluation of seven denoising methods for the QFM signal at SNR levels from $5\mathrm{dB}$ to $5\mathrm{dB}$ under white Gaussian noise: (a) MSE performance of the estimated QFM signals; (b) PSNR performance of the estimated QFM signals.
+Figure 4: Performance evaluation of seven denoising methods for the SFM signal at SNR levels from $5\mathrm{dB}$ to $5\mathrm{dB}$ under white Gaussian noise: (a) MSE performance of the estimated SFM signals; (b) PSNR performance of the estimated SFM signals.
+Additionally, we investigate the denoising performance of the seven methods under conditions of different colored noise.
 The four signals were subjected to various types of colored noise: pink noise, blue noise, and red noise.
-The observation interval is set to $[-5\mathrm{s},5\mathrm{s}]$, and the sampling frequencies are set as follows: $30\mathrm{Hz}$ for example 1, $50\mathrm{Hz}$ for example 2, $150\mathrm{Hz}$ for examples 3 and 4.
+The observation interval is set to $5\mathrm{s},5\mathrm{s}]$, and the sampling frequencies are set as follows: $30\mathrm{Hz}$ for example 1, $50\mathrm{Hz}$ for example 2, $150\mathrm{Hz}$ for examples 3 and 4.
 Table 2 presents the MSE and PSNR values of the output signals through the seven filters under the influence of three types of colored noise.
@@
 For pink and red noise, the Wiener filter performs better than ours for QFM and SFM signals.
+This can be attributed to the spectral characteristics of these signals and the correlation structure of colored noise.
+QFM and SFM signals have rapidly varying or highly oscillatory instantaneous frequencies, and colored noise introduces frequency-dependent distortions that partially align with these variations.
+The Wiener filter, derived from second-order statistics under the minimum MSE criterion, can optimally attenuate noise across the entire spectrum by exploiting such correlations.
+In contrast, the proposed adaptive CCTFD follows the same Wiener filtering principle but performs kernel adaptation in the WVD domain, which may not fully capture the fine-scale structures of rapidly varying signals under colored noise.
+As a result, although the proposed method excels for signals with smoother frequency variations (e.g., LFM and GELFM) and under white or blue noise, its performance can be slightly inferior to the Wiener filter in these particular scenarios.
 However, overall, our method consistently outperforms the fixed kernel function filtering methods and the adaptive radial Gaussian kernel function filtering method.
@@
 MSE (Mean-Square Error)
-Pink noise Margenau-Hill 0.0016-0.0016- 0.0016 0.1592-0.1592- 0.1592 0.0513-0.0513- 0.0513 0.1319-0.1319- 0.1319
-Kirkwood-Rihaczek 0.06760.06760.0676 0.0050-0.0050- 0.0050 0.0330-0.0330- 0.0330 0.1200-0.1200- 0.1200
-Born-Jordan 0.1582-0.1582- 0.1582 0.3053-0.3053- 0.3053 0.1123-0.1123- 0.1123 0.2134-0.2134- 0.2134
-Page 0.09960.09960.0996 0.0360-0.0360- 0.0360 0.0201-0.0201- 0.0201 0.0582-0.0582- 0.0582
-Adaptive radial Gaussian kernel 0.0451-0.0451- 0.0451 0.4237-0.4237- 0.4237 0.0662-0.0662- 0.0662 0.0510-0.0510- 0.0510
-Wiener filter 0.4985-0.4985- 0.4985 0.6911-0.6911- 0.6911 $\mathbf{-0.4699}$ $\mathbf{-0.4268}$
-Adaptive CCTFD $\mathbf{-1.5318}$ $\mathbf{-0.9807}$ 0.3169-0.3169- 0.3169 0.3369-0.3369- 0.3369
-Blue noise Margenau-Hill 0.14370.14370.1437 0.4563-0.4563- 0.4563 0.0637-0.0637- 0.0637 0.3730-0.3730- 0.3730
-Kirkwood-Rihaczek 0.35800.35800.3580 0.3711-0.3711- 0.3711 0.00630.00630.0063 0.2404-0.2404- 0.2404
-Born-Jordan 0.1057-0.1057- 0.1057 0.5265-0.5265- 0.5265 0.2016-0.2016- 0.2016 0.9358-0.9358- 0.9358
-Page 0.32810.32810.3281 0.3355-0.3355- 0.3355 0.02390.02390.0239 0.06820.06820.0682
-Adaptive radial Gaussian kernel 0.0481-0.0481- 0.0481 0.4892-0.4892- 0.4892 0.0671-0.0671- 0.0671 0.0191-0.0191- 0.0191
-Wiener filter 0.4676-0.4676- 0.4676 0.6632-0.6632- 0.6632 0.5208-0.5208- 0.5208 1.0831-1.0831- 1.0831
+Pink noise Margenau-Hill 0.0016-0.0016 0.1592-0.1592 0.0513-0.0513 0.1319-0.1319
+Kirkwood-Rihaczek $0.0676$ 0.0050-0.0050 0.0330-0.0330 0.1200-0.1200
+Born-Jordan 0.1582-0.1582 0.3053-0.3053 0.1123-0.1123 0.2134-0.2134
+Page $0.0996$ 0.0360-0.0360 0.0201-0.0201 0.0582-0.0582
+Adaptive radial Gaussian kernel 0.0451-0.0451 0.4237-0.4237 0.0662-0.0662 0.0510-0.0510
+Wiener filter 0.4985-0.4985 0.6911-0.6911 $\mathbf{-0.4699}$ $\mathbf{-0.4268}$
+Adaptive CCTFD $\mathbf{-1.5318}$ $\mathbf{-0.9807}$ 0.3169-0.3169 0.3369-0.3369
+Blue noise Margenau-Hill $0.1437$ 0.4563-0.4563 0.0637-0.0637 0.3730-0.3730
+Kirkwood-Rihaczek $0.3580$ 0.3711-0.3711 $0.0063$ 0.2404-0.2404
+Born-Jordan 0.1057-0.1057 0.5265-0.5265 0.2016-0.2016 0.9358-0.9358
+Page $0.3281$ 0.3355-0.3355 $0.0239$ $0.0682$
+Adaptive radial Gaussian kernel 0.0481-0.0481 0.4892-0.4892 0.0671-0.0671 0.0191-0.0191
+Wiener filter 0.4676-0.4676 0.6632-0.6632 0.5208-0.5208 1.0831-1.0831
 Adaptive CCTFD $\mathbf{-1.2999}$ $\mathbf{-0.9581}$ $\mathbf{-0.7917}$ $\mathbf{-1.6128}$
-Red noise Margenau-Hill 0.06550.06550.0655 0.3004-0.3004- 0.3004 0.0624-0.0624- 0.0624 0.2233-0.2233- 0.2233
-Kirkwood-Rihaczek 0.12430.12430.1243 0.2567-0.2567- 0.2567 0.0475-0.0475- 0.0475 0.2145-0.2145- 0.2145
-Born-Jordan 0.0907-0.0907- 0.0907 0.3810-0.3810- 0.3810 0.1369-0.1369- 0.1369 0.3623-0.3623- 0.3623
-Page 0.17200.17200.1720 0.2790-0.2790- 0.2790 0.0082-0.0082- 0.0082 0.0851-0.0851- 0.0851
-Adaptive radial Gaussian kernel 0.01050.01050.0105 0.3455-0.3455- 0.3455 0.0743-0.0743- 0.0743 0.0395-0.0395- 0.0395
-Wiener filter 0.8039-0.8039- 0.8039 1.2798-1.2798- 1.2798 $\mathbf{-0.9489}$ $\mathbf{-0.7804}$
-Adaptive CCTFD $\mathbf{-1.7448}$ $\mathbf{-1.3028}$ 0.4032-0.4032- 0.4032 0.3919-0.3919- 0.3919
+Red noise Margenau-Hill $0.0655$ 0.3004-0.3004 0.0624-0.0624 0.2233-0.2233
+Kirkwood-Rihaczek $0.1243$ 0.2567-0.2567 0.0475-0.0475 0.2145-0.2145
+Born-Jordan 0.0907-0.0907 0.3810-0.3810 0.1369-0.1369 0.3623-0.3623
+Page $0.1720$ 0.2790-0.2790 0.0082-0.0082 0.0851-0.0851
+Adaptive radial Gaussian kernel $0.0105$ 0.3455-0.3455 0.0743-0.0743 0.0395-0.0395
+Wiener filter 0.8039-0.8039 1.2798-1.2798 $\mathbf{-0.9489}$ $\mathbf{-0.7804}$
+Adaptive CCTFD $\mathbf{-1.7448}$ $\mathbf{-1.3028}$ 0.4032-0.4032 0.3919-0.3919
 PSNR (Peak Signal-to-Noise Ratio)
-Pink noise Margenau-Hill 3.03233.03233.0323 5.27665.27665.2766 3.52373.52373.5237 4.33064.33064.3306
-Kirkwood-Rihaczek 2.34002.34002.3400 3.74103.74103.7410 3.34313.34313.3431 4.21104.21104.2110
-Born-Jordan 4.61904.61904.6190 7.07167.07167.0716 4.13394.13394.1339 5.14945.14945.1494
-Page 2.04332.04332.0433 3.73703.73703.7370 3.21093.21093.2109 3.59633.59633.5963
-Adaptive radial Gaussian kernel 3.47583.47583.4758 7.31207.31207.3120 3.67363.67363.6736 3.52163.52163.5216
-Wiener filter 8.83148.83148.8314 10.319210.319210.3192 $\mathbf{8.0556}$ $\mathbf{7.8716}$
-Adaptive CCTFD $\mathbf{18.5390}$ $\mathbf{12.8503}$ 6.19786.19786.1978 6.57996.57996.5799
-Blue noise Margenau-Hill 1.58651.58651.5865 7.62287.62287.6228 3.64683.64683.6468 7.23777.23777.2377
-Kirkwood-Rihaczek 0.5454-0.5454- 0.5454 6.81086.81086.8108 2.94692.94692.9469 5.42785.42785.4278
-Born-Jordan 4.15574.15574.1557 8.37578.37578.3757 5.03425.03425.0342 13.017913.017913.0179
-Page 0.2493-0.2493- 0.2493 6.47646.47646.4764 2.77082.77082.7708 2.32972.32972.3297
-Adaptive radial Gaussian kernel 3.50463.50463.5046 7.90757.90757.9075 3.68133.68133.6813 3.20163.20163.2016
-Wiener filter 8.10828.10828.1082 9.89379.89379.8937 8.68998.68998.6899 14.400014.400014.4000
+Pink noise Margenau-Hill $3.0323$ $5.2766$ $3.5237$ $4.3306$
+Kirkwood-Rihaczek $2.3400$ $3.7410$ $3.3431$ $4.2110$
+Born-Jordan $4.6190$ $7.0716$ $4.1339$ $5.1494$
+Page $2.0433$ $3.7370$ $3.2109$ $3.5963$
+Adaptive radial Gaussian kernel $3.4758$ $7.3120$ $3.6736$ $3.5216$
+Wiener filter $8.8314$ $10.3192$ $\mathbf{8.0556}$ $\mathbf{7.8716}$
+Adaptive CCTFD $\mathbf{18.5390}$ $\mathbf{12.8503}$ $6.1978$ $6.5799$
+Blue noise Margenau-Hill $1.5865$ $7.6228$ $3.6468$ $7.2377$
+Kirkwood-Rihaczek 0.5454-0.5454 $6.8108$ $2.9469$ $5.4278$
+Born-Jordan $4.1557$ $8.3757$ $5.0342$ $13.0179$
+Page 0.2493-0.2493 $6.4764$ $2.7708$ $2.3297$
+Adaptive radial Gaussian kernel $3.5046$ $7.9075$ $3.6813$ $3.2016$
+Wiener filter $8.1082$ $9.8937$ $8.6899$ $14.4000$
 Adaptive CCTFD $\mathbf{16.0114}$ $\mathbf{12.6000}$ $\mathbf{10.9354}$ $\mathbf{19.6753}$
-Red noise Margenau-Hill 2.38882.38882.3888 6.39986.39986.3998 3.63603.63603.6360 5.25625.25625.2562
-Kirkwood-Rihaczek 1.77741.77741.7774 5.94565.94565.9456 3.48813.48813.4881 5.15595.15595.1559
-Born-Jordan 4.01114.01114.0111 7.41677.41677.4167 4.38054.38054.3805 6.63486.63486.6348
-Page 1.36061.36061.3606 5.98635.98635.9863 3.09253.09253.0925 3.86283.86283.8628
-Adaptive radial Gaussian kernel 2.99442.99442.9944 6.62676.62676.6267 3.75523.75523.7552 3.40613.40613.4061
-Wiener filter 11.883211.883211.8832 $\mathbf{16.7052}$ $\mathbf{13.1156}$ $\mathbf{11.9071}$
-Adaptive CCTFD $\mathbf{20.4654}$ 16.149516.149516.1495 7.04967.04967.0496 7.20567.20567.2056
+Red noise Margenau-Hill $2.3888$ $6.3998$ $3.6360$ $5.2562$
+Kirkwood-Rihaczek $1.7774$ $5.9456$ $3.4881$ $5.1559$
+Born-Jordan $4.0111$ $7.4167$ $4.3805$ $6.6348$
+Page $1.3606$ $5.9863$ $3.0925$ $3.8628$
+Adaptive radial Gaussian kernel $2.9944$ $6.6267$ $3.7552$ $3.4061$
+Wiener filter $11.8832$ $\mathbf{16.7052}$ $\mathbf{13.1156}$ $\mathbf{11.9071}$
+Adaptive CCTFD $\mathbf{20.4654}$ $16.1495$ $7.0496$ $7.2056$
+To further validate the practical effectiveness of the proposed adaptive CCTFD, we conduct experiments on real-world data.
+Specifically, the RadChar dataset is used, focusing on the RadChar-Tiny subset.
+A segment of the first $7500$ samples is extracted and contaminated with additive white Gaussian noise under three SNR levels: $1\mathrm{dB}$, $2\mathrm{dB}$, and $3\mathrm{dB}$.
+The MSE and PSNR results are summarized in Table 3.
+As observed, the proposed adaptive CCTFD consistently achieves the lowest MSE and highest PSNR across all SNR levels, demonstrating superior denoising performance.
+These results indicate that the proposed adaptive CCTFD is effective not only for simulated signals but also for real radar data, highlighting its practical applicability and robustness.
+Table 3: The MSE and PSNR of seven filtering methods for the RadChar dataset under white Gaussian noise.
+Noise level (SNR) 3-3 2-2 1-1
+MSE Margenau-Hill $0.0639$ $0.0567$ $0.0493$
+Kirkwood-Rihaczek $0.0640$ $0.0574$ $0.0509$
+Born-Jordan $0.0491$ $0.0406$ $0.0319$
+Page $0.1006$ $0.0965$ $0.0924$
+Adaptive radial Gaussian kernel $0.1176$ $0.0983$ $0.0955$
+Wiener filter 0.0169-0.0169 0.0407-0.0407 0.0672-0.0672
+Adaptive CCTFD $\mathbf{-0.0317}$ $\mathbf{-0.0500}$ $\mathbf{-0.0675}$
+PSNR Margenau-Hill $3.7697$ $3.8163$ $3.8651$
+Kirkwood-Rihaczek $3.5877$ $3.6311$ $3.6764$
+Born-Jordan $3.8719$ $3.9283$ $3.9876$
+Page $3.5575$ $3.5916$ $3.6258$
+Adaptive radial Gaussian kernel $3.6309$ $3.7486$ $3.7686$
+Wiener filter $3.8878$ $4.0373$ $4.2206$
+Adaptive CCTFD $\mathbf{4.1225}$ $\mathbf{4.2956}$ $\mathbf{4.4592}$
 5 Conclusion
@@
 Under white Gaussian noise, our method outperforms the ordinary Wiener filter, while under colored noise, it performs comparably to the Wiener filter.
-[…]
-Acknowledgements This work was supported by the National Natural Science Foundation of China (Grant No 61901223), the Jiangsu Planned Projects for Postdoctoral Research Funds (Grant No 2021K205B), the Foundation of Key Laboratory of System Control and Information Processing, Ministry of Education (Grant No Scip20240121), and the Foundation of Key Laboratory of Computational Science and Application of Hainan Province (Grant No JSKX202401).
+Experiments on real-world radar data from the RadChar dataset further confirm that the proposed method achieves superior MSE and PSNR, demonstrating its practical applicability and robustness.
 Appendix A Proof of Eq. (10)
@@
 […]
+[…]
+Author contributions M.J.C.
+conceived the study, developed the methodology, performed the experiments, and drafted the manuscript.
+Z.Z.C.
+supervised the research, contributed to the theoretical analysis, and provided critical revisions to the manuscript.
+All authors reviewed and approved the final version of the manuscript.
+[…]
+Funding This research is supported by the Open Foundation of Hubei Key Laboratory of Applied Mathematics (Hubei University) (HBAM202404), and the Key Laboratory of System Control and Information Processing, Ministry of Education (Scip20240121).
+[…]
+Data availability The experiment data used to support the findings of this study are available from the corresponding author upon request.
+Declarations
+[…]
+Competing interests The authors declare no competing interests.